### v1

# AI-Powered Urban Transportation Digital Twin: Methods and Applications

Abstract
We present a survey paper on methods and applications of digital twins (DT) for urban traffic management. While the majority of studies on the DT focus on its “eyes,” which is the emerging sensing and perception like object detection and tracking, what really distinguishes the DT from a traditional simulator lies in its “brain,” the prediction and decision making capabilities of extracting patterns and making informed decisions from what has been seen and perceived. In order to add values to urban transportation management, DTs need to be powered by artificial intelligence and complement with low-latency high-bandwidth sensing and networking technologies. We will first review the DT pipeline leveraging cyberphysical systems and propose our DT architecture deployed on a real-world testbed in New York City. This survey paper can be a pointer to help researchers and practitioners identify challenges and opportunities for the development of DTs; a bridge to initiate conversations across disciplines; and a road map to exploiting potentials of DTs for diverse urban transportation applications.
footnotetext: Xuan Di, Yongjie Fu, Mehmet K.Turkcan, and Zhaobin Mo are with the Department of Civil Engineering and Engineering Mechanics, Columbia University. Emails: {sharon.di, yf2578, mkt2126, zm2302}@columbia.edufootnotetext: Xuan Di, Zoran Kostic, and Gil Zussman are also with the Data Science Institute, Columbia University.footnotetext: Mahshid Ghasemi, Chengbo Zang, Abhishek Adhikari, Zoran Kostic, and Gil Zussman are with the Department of Electrical Engineering, Columbia University.Emails: {mg4089, cz2678, abhishek.adhikari, zk2172, gil.zussman}@columbia.edu
Keywords: Digital twin, AI, Urban transportation management, Survey
1 Introduction
Urban transportation systems are complex to model and simulate, due to heterogeneous road users (such as cars, pedestrians, cyclists, scooters) interacting in multimodal traffic environments consisting of public and private travel modes. With fast-changing traffic evolution in time and space, traffic simulation if improperly calibrated, might produce traffic management strategies that largely deviate from the reality, potentially leading to suboptimal or even detrimental outcomes. With ubiquitous sensors in smart cities, it is the time to augment conventional traffic simulators, many of which were developed in an era when only “small data” became available. Emerging traffic sensors are expected to generate big volumes of data, transmitted via communication networks and processed on edge cloud computing with artificial intelligence (AI) for real-time traffic management. Such a transformation poses challenges that call for the development of a new paradigm, namely, digital twin (DT), which will push the envelope in urban transportation management.
Literally, DT is the digital replica of a physical object or asset [48], where a digital world mirrors a physical world for real-time diagnosis, prognosis, and decision making. Recent years have seen a growing amount of studies on DT in various domains, ranging from manufacturing [48], aerospace [76], to smart cities [82]. In particular, there is a sizable body of articles on transportation DTs powered by vehicular technologies such as connectivity and autonomy [134, 162, 57, 199, 80]. With such a growing trend, we would like to restrict the scope of this survey paper. Here we are particularly interested in urban traffic settings where cars have to interact with vulnerable road users (VRU) (i.e., non-motorists such as pedestrian, bicyclists, other cyclists, or persons on personal conveyance [153]), who are an integral part and also potential users of the DT. We will primarily focus on two use cases, namely, intersection safety warning and network wide traffic signal control, for improved traffic safety and efficiency. The literature on DTs for urban traffic management is lacking, partly because the development of a DT for a system is non-trivial, particularly involved with human participants like VRUs. The development of such a DT must engage with expertise in sensing, communication, computing, and human-centric perspectives. Accordingly, we strongly believe that a comprehensive review is warranted on trans-disciplinary topics.
This paper presents a survey paper on methods and applications of DT for urban traffic management. We will depict a DT pipeline prototype, leveraging the architecture of cyberphysical systems and AI methods. We will propose a DT for real-time traffic monitoring and optimization [127], based on an existing physical testbed deployed in New York City (NYC) leveraging cutting-edge sensing, communication, computing, and AI-based automation. The overall contributions of this paper include: (1) introducing AI methods used in the DT pipeline; (2) exploring the architecture of transportation DTs and propose a prototype for reference; and (3) identifying gaps and directions.
The rest of this paper is organized below. Section 2 unifies various concepts of DT and introduces the relationship between cyberphysical transportation systems and DTs; Section 3 reviews the literature on sensing and communication technologies, object detection and tracking, and urban transportation use cases. Section 4 demonstrates the architecture of our proposed DT building on a real-world testbed. Section 5 concludes our work and presents potential research directions as well as open questions.
2 Preliminaries
2.1 Variants of DT definitions
The definition of DT has evolved rapidly. Despite presenting its own version, these articles all share common elements and resemble certain characteristics [48, 114]. In general, there is a physical world (aka. the physical) and a digital world (aka. the digital). The physical world evolves in time and space. To ensure that the physical system is run in a desired direction, it requires close monitoring, operation, and management. Thus, the role a digital world plays is to model and simulate the dynamics of the physical world in a synchronized fashion, so that the digital can also predict the future states of the physical precisely, which offers a ground for optimal decision making. The physical and the digital exchange data and information flows via a two-way communication. Specifically, the physical sends the data of its own state to the digital, and the digital feeds back the actuation signals to the physical. The actuation would trigger a change in the stage of the physical, and the updated state is sensed and sent back to the digital again. This iterative process runs between the physical and the digital as time unfolds. The sequential states of the physical should move towards a more desired state than without a DT. In other words, the ultimate goal of a DT development is to add values to the physical for improved safety and efficiency.
Depending on whether there is any interaction between the physical and the digital via data and/or information exchange, there are three types of digital twinning concepts [47], namely,
  1. 1.
    digital model: with no automatic data exchange;
  2. 2.
    digital shadow: with only one-way data exchange from the physical to the digital;
  3. 3.
    digital twin: with two-way communication.
Let us specify a definition of DT in transportation.
Definition 2.1
Traffic digital twin is a digital system integrating the pipeline from object detection and tracking, resource allocation, edge-cloud computing and communication, for online simulation, operation, control and management. It is updated online using continuously fed data collected from the physical and send control policies or issue warnings back to the physical, leveraging big data and AI tools. Traffic DTs are closed-loop with two-way communication, where data, information, and control signals are exchanged with the physical sequentially and reiteratively.
2.2 Cyberphysical transportation systems
Definition 2.2
Cyberphyscial systems (CPS) [49] are smart systems that include engineered interacting networks of physical and computational components. CPS holds great potential to enable real-time applications thanks to emerging technologies in sensing, communication, and computing.
2.2.1 CPS as DT technological enabler
A transportation CPS interlinks physical and cyber layers, where the cyber layer consists of sensing, networking, computing, and traffic management application modules (see Fig. 1). The DT encloses the cyber layer, and relies on all the modules for two-way interaction with the physical.
To enable the technological development of a DT, a physical testbed is needed as a platform for sensing, computing, experimentation, evaluation, as well as design constraints determination. In Secs. 3.1-3.3, we will review the CPS technological enablers needed for the development of a DT, and examine the testbed used for our proposed DT in Sec. 4.1.1.
2.2.2 DT as CPS testing tool
An automatic system like CPS demands high degree of autonomy for urban traffic management, with increased complexity and scale. Thus, DT becomes crucial for the test and verification of a CPS. We need to verify the system behavior of a CPS to ensure that it performs as desired [141], such as safety, efficiency, timeliness, and accuracy.
Due to the integration of cyber and physical components, as well as their two-way coupling via communication and networking, testing a CPS is non-trivial. Given that the digital is expected to be the twin of the physical, it is natural to compare the difference in the outputs of both the physical and the DT using “Grieves performance test” [48]. Hardware-in-the-loop testing, such as vehicle-in-the-loop [172], could help test directly the performance of some physical components “live,” while allowing the rest to be simulated within a DT. For DTs that involve the feedback to humans, human-in-the-loop test is widely used on driving simulators [183, 79]. Recent years have seen a growing trend of using augmented reality and virtual reality (AR/VR) or Metaverse to engage humans as pedestrians, cyclists, or scooters in virtual environments without inducing real-world risks [164, 72].
Since there does not exist a unified approach for the DT testing in the context of transportation, in Sec. 3.4, we will review evaluation metrics employed by various use cases.
3 Related work
In this section, we will review technological enablers of DT along the pipeline depicted in Fig. 1.
Figure 1: Pipeline of a CPS-structured DT.
3.1 Sensing and perception
Sensors are the “eyes” (and “ears”) of a DT. Tab. 2 summarizes the pros and cons of each sensing technology, namely, mobile devices, on-board vehicles, and roadside infrastructure, for urban traffic applications. The former two are mobile sensors with wider spatial coverage but challenge in precision because of moving references, while the latter at fixed locations could face limited sensing ranges and coverage.
Table 1: Pros and cons of sensors for urban traffic applications (partially adapted from [154, 194])
Sensor Purpose Advantages Disadvantages
Mobile GPS Pedestrian localization 1. Offers global coverage and compatibility with a wide range of devices.
2. Offers reliability and easy access, as it is widely adopted in consumer devices.
1. Accuracy within a few meters, which may be insufficient for safety-critical applications.
2. Poor signals through obstacles.
3. A low update rate for real-time safety applications.
UWB Pedestrian localization 1. Achieves high accuracy, often within a few centimeters, making it suitable for precise indoor and short-range outdoor applications.
2. Operates with low latency, providing real-time updates of position data.
1. Requires anchors to be installed at each corner of the designated area before localization.
2. Operates in a frequency range that may overlap with other wireless technologies, such as Wi-Fi, Bluetooth, or cellular networks [26].
On-board Vehicle Radar Obstacle Detection 1. Outperforms other sensor types at far distances.
2. Detects vehicle speed and position accurately without the need for calibration.
3. Protects privacy, as this sensor type does not record identifiable images of road users.
1. Performs best only when objects move toward or away from the sensor.
2. A Limited number of classes that can be identified due to the lack of color and resolution.
3. Limited field of view when the range is far [124].
Camera Obstacle and lane detection 1. Maintains good resolution when the FOV is wide.
2. Has a long horizon.
1. Difficulties in measuring speed and distance.
2. Performs poorly in bad weather conditions.
Lidar Obstacle detection, 3D mapping 1. Measures distance accurately.
2. Constructs 3D models robustly.
3. Shows promising performance in poor weather.
1. Detects nearby objects poorly [124].
2. Demands high data processing requirements.
3. A shorter effective range than radar.
Accelero -meter Acceleration, driving behavior 1. Detects braking, turning, and accelerating accurately.
2. Can be integrated with the vehicle system for behavior analysis [98].
1. Performs poorly for slow or subtle movements.
2. Detects poorly in the presence of noise.
Roadside Infrastructure Camera Object detection at intersections 1. Provides more details, compared to radar and LiDAR, that can be used to differentiate types of vulnerable road users.
2. Covers a larger area where pedestrians are not confined to a narrow path, such as when people crossing midblock [150].
1. Performs poorly in adverse weather conditions.
2. Difficulty in long-term use due to the cost, power supply, and quantity.
3. No privacy protection.
Acoustic Ultrasonic Lane occupancy/vehicle speed 1. Collects data on multiple lanes.
2. Operates during both day and night.
1. Undercounts or overestimates speed.
2. Performs poorly in severe weather.
MMWave Radar Vehicle localization, speed measurement 1. Features a compact size and is easy to install.
2. Offers low latency, within 30ms [161].
3. Penetrates through non-metallic objects.
1. Provides low angular resolution.
2. Measures elevation poorly.
3. Has difficulty with real-time calibration [186, 187].
3.2 Object detection and tracking
Object detection identifies and classifies objects within the environment using sensors like cameras, Lidar, and Radar. Object tracking is the process of monitoring the detected objects over time to determine their position and movement. Multi-object tracking is concerned with maintaining the identity of the objects and generating their trajectories. Trajectory prediction involves forecasting the future paths of detected and tracked objects. These tasks highly rely on training datasets for urban traffic scenes. Note that there is a much larger size of public datasets collected from on-board vehicles [31], but fewer from other sensor types. Here we thus summarize commonly used and emerging datasets from non-vehicle sensors in Tab. 1.
Table 2: Public datasets for urban traffic scenes (On-board vehicle datasets can be found in [31].)
Sensor location Dataset Purpose Sensor Setup Collection region
Infrastructure VIRAT [117], Constellation [152] Urban object detection and tracking, visual event recognition RGB cameras Public outdoor spaces in China, city intersection in New York
Rope3D [181] 2D/3D Road-side object detection, multi-view RGB cameras, LiDAR Streets in Beijing
DeepSense 6G [11] 2D/3D object detection, sensor fusion RGB cameras, mmWave Phase Arrays, LiDAR, Radar Various indoor and outdoor spaces
Aerial VisDrone [202], NGSim [116], highD [67], rounD [68], DOTA [171], CitySim [198] Aerial object detection and tracking, trajectory forecasting Drone-based RGB cameras Urban spaces in China and Aacheny, intersections in California and Florida
Misc CoCo [85], ADE20K [200], Cityscapes [200] Object detection, semantic segmentation RGB cameras Various indoor and outdoor spaces
Object detection has been studied extensively for urban applications. Single-stage object detectors, following the original single shot multibox detector [93] and You Look Only Once (YOLO) [126] architectures, have become popular due to their real-time deployment capabilities. In the last few years, transformer-based object detection approaches, competitive with YOLO models, have emerged as the state-of-the-art in object detection when designed to be deployed in real-time settings [197]. Recent progress in YOLO object detection performance has been enabled through multiple small tricks in architecture and training that all together provide significant improvements in empirical performance [155].
When multiple camera views are available, 3D object detection has been studied heavily for autonomous driving [46, 146], as well as for infrastructure-based 3D object detection [117, 152, 181, 11]. Many approaches to 3D object detection use object queries [160], bird’s-eye view transformations [173], or a combination of the two [90].
To build models that make weaker assumptions regarding the sensors, some models have considered the harder task of monocular 3D object detection. MonoCon uses extra regression head branches for learning auxiliary contexts, that are then discarded during inference [95]. DEVIANT is a model architecture equivariant to depth translations [70]. MonoLSS introduces a learnable sample selection module to improve the stability and reliability of the model at test time [81]. Different models have been proposed for infrastructure-based 3D object detection, as many models developed for vehicle-side perception make strong assumptions regarding the position of the cameras. BEVHeight predicts height to the ground to support 3D object detection [175]. CoBEV combines depth and height features to further improve the performance of infrastructure-based 3D object detection [137]. MonoUNI presents the idea of normalized depth, which makes depth prediction independent of camera pitch angle and focal length [61].
To improve the limitations of camera-only perception methods, different sensor combinations are explored. For example, LiDARs or radars, combined with cameras, can detect objects in scenarios where using only the camera is insufficient, such as extreme lighting and weather conditions, or anomalous situations where the camera data is significantly out-of-distribution. Sensor fusion for self-driving cars is now being studied including sensor data for these modalities [23, 87]. These methods often involve the projection of camera, radar and LiDAR features independently to a bird’s-eye view feature space, wherein an aggregation function could be used to merge the features extracted from these different sensors.
Multi-object tracking involves matching newly detected objects with the existing ones by their inter-frame positional and visual similarity information [18, 167]. ByteTrack improves the traditional Hungarian-algorithm-based matching paradigm to gather more comprehensive information [193]. BoT-SORT further incorporates advanced object re-identification modules and a refined Kalman filter for more accurate performances [5]. BoostTrack explores novel distance and shape similarity measurements to deal with ambiguity caused by unreliable detection results [142, 143]. Modern MOT pipelines are often implemented directly downstream to object detection in the tracking-by-detection paradigm.
Majority of modern architectures for predicting future trajectories of detected objects adopt Recurrent Neural Networks which is responsible for predicting future object positions based on their historical coordinates, together with generative components which handles the variation and flexibility in social interactions [8, 54]. Neural Social Physics model [185] incorporates learnable parameters into explicit physics models built on top of neural networks. SemanticFormer [147] seeks more structural and humanized environment understanding by constructing a semantic knowledge graph. TrajNet++ [66], TDOR [53], and CASPNet++ [135] predict the distributions of future trajectories based on occupancy grid maps. Models like FRM [121] and PPT [86] decompose the prediction task by taking a multi-stage approach. Larger amounts of data comprised of multiple modalities and more comprehensive frameworks have shown increasing importance as is demonstrated by UniTraj++ [39]. Unlike the tracking algorithms, specific training or fine-tuning is often required before the deployment of trajectory forecasting models to unseen scenarios.
3.3 Real-time video analytics
Developing end-to-end real-time video analytics systems on a large scale for time-sensitive and safety-critical traffic and crowd management applications presents challenges. Video analytics requires the collection and processing of large volumes of video data, which can be resource-intensive and costly. Optimizing computation and network resource usage while maintaining or enhancing the accuracy of analytical results can be challenging. This challenge is further complicated by the need to adapt to varying network conditions, computational resources, and dynamic scene changes in real-time. Tab. 3 provides a comparative analysis of various approaches to address these challenges. The approaches differ in their focus–some prioritize reducing latency and resource consumption, while others emphasize maintaining or enhancing accuracy, especially under constrained conditions. This comparison provided in Tab. 3 highlights the trade-offs inherent in real-time video analytics and emphasizes different optimization methods to balance throughput, accuracy, energy consumption, and computational efficiency across diverse deployment scenarios, including edge devices, cloud platforms, and hybrid environments.
Table 3: Comparison of various approaches and optimization objectives in video analytics.
References Approach Optimization objective
SPINN [73], Adaptive offloading [190], Shoggoth [158], Sniper [94], JAVP [176], Auction-base [41] Distributed DNN inference over end devices, edge, and cloud. Optimize throughput, accuracy, and energy consumption under varying network conditions.
CEVAS [174], SAHI [7], CrossRoI [51], Elf [192] Adaptive RoI assignment and frame sampling. Reduce the bandwidth consumption and enhance accuracy.
AdaMask [91], Respire [28], CrossVision [189], VaBUS [156] Leveraging redundant regions on frames and background understanding. Minimize network and computation overhead while ensuring high accuracy.
Elf [192], Mobile edge analytics [84], Sniper [94], JAVP [176], Auction-base [41] Video analytics query scheduling and resource allocation over multiple edge devices. Reduce latency and increase computation resource utilization.
CEVAS [174], Edge-assisted serverless [163] Adaptive model selection. Enhance performance with limited computation resources.
Shoggoth [158], Edge-assisted [65] Online model fine-tuning and model switching. Improve the accuracy and efficiency of real-time video inference on edge devices in changing video scenes.
DAO [112], VaBUS [156], AccMPEG [34], AdaMask [91], ILCAS [169] Adaptive video encoding and compression parameters. Balance low latency, high accuracy, and low compute overhead on edge devices.
AdaDSR [20], AccDecoder [184] Camera-side downsampling and server-side super-resolution upsampling. Balance the trade-offs among accuracy, network cost, and computational cost.
MadEye [168], WiseCam [60] Dynamical orientation adaption of pan-tilt-zoom (PTZ) cameras. Boost the overall accuracy while maintaining the resource cost.
EAIS [179], EALI [178], SERAS [56] Use of an energy-aware scheduler that effectively coordinates batching and dynamic voltage frequency scaling (DVFS) settings. Minimize energy consumption for CNN inference services on high-performance GPUs while meeting latency of Service-Level Objectives.
3.4 Use cases
The value of a DT for urban transportation management ranges from traffic state prediction, spatiotemporal traffic flow forecasting, to urban planning and policy making. Here we focus on two that are particularly important in urban settings.
3.4.1 VRU safety warning: Safety-critical application
Safety-critical applications, making automated life-and-death decisions such as collision avoidance warning between automobiles and pedestrians, need to activate at the precise time and the right moment with bounded latency. Safety-critical systems are often time-critical and reactive, because they need to react to external signals in precise time [69], and require tasks to be executed in a timely manner. Theses systems hold the substantial potential to enhance road safety and save lives. They are, however, risky to test and run, because rare events like automobile collisions are challenging and unethical to replicate. Thanks to emerging technologies in ubiquitous sensing, low-latency high-bandwidth communication, high-speed computing and AI, safety critical applications could potentially be modeled, simulated, processed, and tested in a DT. The applications of DTs on safety-critical scenarios, however, remains understudied, because it necessities short runtime and real-time reaction, posing high requirements for communication and computing technologies, pipeline architecture design, and testing.
We summarize the safety-critical applications to address conflict risks between vehicles and VRUs in Tab. 4, and offer an outlook of safety guarantee related methods in Sec. 5.1.4.
Table 4: Literature review on safety warning to VRUs.
Ref Technology Objective Risk Assessment Metrics Evaluation Method/Metric
[118] Sensing: roadside sensors and VRU smartphones
Communication&Networking: 5G, MNO infrastructure, and ITS-G5
Computing: Edge-cloud hybrid computing
Reduces latency and optimizes resource utilization through dynamic service placement. The distance between the VRU and vehicles End-to-end delay: 200 ms
[36] Sensing: Real-time camera detection, YOLOv7
Communication&Networking: 5G, 802.11p, C-V2X, Vehicular Basic Safety Message (BSM)
Computing: Local server equipped with GPU
Develops a video-based vehicular BSMs method with lower error and latency that outperforms the cellular vehicle-to-everything (C-V2X) method. N/A 1. End-to-end delay: <100absent100<100< 100  ms
2. Localization/Speed accuracy
[113] Sensing: GPS on smartphones
Communication&Networking: LTE, Node B, Wi-Fi, Cooperative Awareness Message (CAM)
Computing: CAM server deployed at edge/cloud
Proposes a system using commercial devices and standard messages for road user communication. The distance between VRUs and nearby entities Latency from VRU to CAM <50absent50<50< 50 ms
[99] Sensing: Camera, Android phone’s GPS module
Communication&Networking: 3G, 4G, 5G, and MQTT protocol
Computing: Coral Edge TPU with TensorFlow lite
Develops a traffic safety system using edge computing and 5G to deliver low-latency warnings. The coordinates of pedestrians and cyclists in a driver’s blind spots Latency:
1. 4G - 109.35 ms
2. 5G - 90.95 ms
[83] Sensing: N/A
Communication&Networking: Wi-Fi, C-V2X, 802.11p, ITS-G5, Cooperative Awareness Message
Computing: Edge computing server, smartphones
[…]
2. MAC channel utilization
[148] Sensing: Cameras, Radar, YOLOv3 on NVIDIA JETSON, On Board Units (OBUs)
Communication&Networking: 5G, Fiber, LTE, ITS-G5, and MQTT
Computing: Road side units
Develops a system with sensing and communication, along with fusion and collision detection algorithms, to predict potential collisions and warn VRUs. The distance between VRUs and nearby entities 1. End-to-end latency <300absent300<300< 300 ms
2. Distance error of vehicles and VRUs
[43] Sensing: Camera, real time object detection
Communication: LTE, Wi-Fi, and MQTT protocol
Computing: GPU on server end
Develops a real-time system with a mobile application to warn pedestrians to avoid vehicle and walker collisions. Time to collision (TTC) 1. End-to-end latency 400 ms.
2. Simulation.
[164] Sensing: CAN, GNSS, roadside, cameras, Google MediaPipe Posekeypoint
Communication&Networking: Fiber, and Simulated V2X
Computing: Server end GPU and portable GPU
Develops a DT framework for connected vehicles and pedestrian in-the-loop simulation. Test it with a V2P collision warning use case. TTC 1. Speed
2. Brake point
3. Distance to conflict point
Ours Sensing: Camera, YOLOv8 object detection.
Communication&Networking: Fiber, LTE, and MQTT protocol
Computing: Server end GPU
Develops a DT pipeline to demonstrate use cases (including intersection safety warning and ATSC) in urban settings. TTC 1. Accuracy for the warning issued
2. Granular latency per frame
3.4.2 Networkwide adaptive traffic signal control
Learning methods for network wide adaptive traffic signal control (ATSC), including reinforcement learning (RL) and federated learning (FL), aim to learn optimal traffic signal control strategies in traffic environments with unknown state transitions and unknown rewards. Assuming that traffic signal controllers are intelligent autonomous agents, their goal is to make sequential decisions to optimize an objective function, such as minimization of traffic delay or maximization of traffic throughput. To train an optimal policy for traffic signal controllers, a traffic simulator or DT, needs to be built, where dynamic traffic is simulated in the training environment aligned with the testing environment in principle. The trained policies leveraging such a DT need to be executed in the real world, which triggers updated environments for future decision making. Related work are summarized in Tab. 5.
Table 5: Summary of papers using (F)RL for ATSC
Ref. RL FL DT Simulator Real-world Demand Real-world Network
[63] DDPG SUMO
[30] SUMO
[71] A2C Vissim
[64] Vissim
[201] SUMO
[42] A3C FedAvg SUMO
[182] DQN FedAvg Cityflow
[58] PPO FedAvg SUMO
[16] DQN FedAvg SUMO
[109] A2C SUMO
[77] A2C SUMO
[27] A2C SUMO
[52] PPO SUMO
FedAvg: Federated Averaging, DDPG: Deep Deterministic Policy Gradient, A2C: Advantage Actor Critic, A3C: Asynchronous Advantage Actor Critic, PPO: Proximal Policy Optimization, DQN: Deep Q Learning.
3.5 Close the loop
To realize the value of a DT, the feedback from the digital to the physical is the key. The type of feedback and to what entities would define the technological constraints.
  • Feedback to VRUs (super local): When users react to an issued warning, the state of the environment for the VRU would change, and the DT could update its status about whether continuous warning is needed. Normally the warning is safety related, thus demanding low latency and high bandwidth.
  • Feedback to cars (super local): The feedback to cars includes hierarchies: (1) strategic driving feedback, such as routing recommendations, allowing higher latency and lower bandwidth; (2) tactical feedback on tactical driving decisions, such as lane-change, with medium latency and bandwidth; and (3) operational feedback such as throttle control, demanding (ultra) low latency and bandwidth. When cars change their driving maneuvers, the traffic environment needs to be updated and unsafe outcomes could be potentially avoided.
  • Feedback to traffic operators and infrastructure (local or networkwide): For traffic operators, DTs could provide anomaly detection for diagnosis, and predict future traffic states or collisions for prognosis, which would help with variable speed limit, dynamic speed advisory, lane usage adaptation, ramp metering, traffic signal timing optimization, and network wide traffic management, with increasing requirements for latency and bandwidth.
  • Feedback to city planners (citywide): In the long-term, a DT could provide aggregate traffic information to city planners to assess the potential societal impact of new infrastructure planning. Thus, the technological requirement is relatively low, but the collected data needs to cover a relatively large spatiotemporal range for citywide decision making,
3.6 Communication and networking
A DT for safety-critical applications requires real time communications with aggressively low latency. We explore issues related to low latency communications, and survey component technologies and protocols that can be utilized to achieve very low latency.
3.6.1 Real-time requirements and low latency targets
Sensor and control data in a real time system is subject to latency created by the stages, namely, (1) data acquisition from traffic participants (such as camera recordings and encoding, harvesting data from autonomous vehicles, and collect information from fiber); (2) transmission of data across communications links from sensors to inferencing servers using communications protocols such as Transmission Control Protocol (TCP), Unreliable Data Protocol (UDP) and Real-Time Streaming Protocol (RTSP); (3) data preprocessing (video decoding, and cropping); (4) AI inferencing; (5) higher level reasoning about required feedback to traffic participants; and (6) sending feedback to traffic participants across communications links via low-latency broadcast or dedicated channels over wired and wireless.
Smart city applications can be grouped according to their latency requirements. Many, if not all, pedestrian-associated application (facilitated by pedestrian detection/observations and message notifications) are likely to expect the round trip delay in the range of a couple of seconds. Such latency can be supported by contemporary cameras, communication protocols and inferencing engines. Applications which would attempt to close the observation/notification loop for vehicles moving at about 10kmph may expect latency in tens of millisecond. Using conventional video compression, RTP/RTSP streaming and edge computing is inadequate to support such latency. This presents the opportunity to pursue novel engineering solutions and research problems.
3.6.2 Communication techniques and protocols
Ultra-Reliable Low Latency Communications (URLLC), a key component of 5G wireless, can help achieve the low latency targets. Along with enhanced mobile broadband (eMBB) and massive machine-type communication (mMTC), URLLC [123] represents one of the three main capabilities of 5G New Radio (5G NR), as standardized by the 3rd Generation Partnership Project (3GPP). In the context of transportation systems, URLLC aims to deliver up to 99.999% reliability and single-digit millisecond latency [2]. However, meeting these performance metrics is challenging in practice due to complex channel environments, particularly in dense urban areas, which can reduce reliability. For intelligent transportation systems, where URLLC may be used as infrastructure backhaul, the target is an end-to-end latency of 30 ms [1].
An emerging technology in this space is Cellular-Vehicle-to-Everything (C-V2X), which has largely replaced the earlier Wi-Fi-based Dedicated Short-Range Communications (DSRC). Unlike DSRC, C-V2X leverages cellular networks, allowing network providers to offer always-on connectivity, which is a critical feature for time-sensitive applications. Additionally, private 5G networks are being developed to ensure this level of connectivity, overcoming the congestion and range limitations inherent to Wi-Fi [37]. In transportation systems, active collaboration between wireless service providers and vehicle manufacturers is in progress to integrate private 5G networks into vehicular networks [3].
Message Queuing Telemetry Transport (MQTT) is a lightweight communication protocol that offers efficient, reliable, and scalable data transfer necessary for real-time, safety-critical applications. Such a protocol can efficiently collects and distributes traffic data from numerous sensors to update DTs in real-time, and facilitates fast communication of incidents and emergencies, allowing timely interventions. Key features include [122]:
  • Lightweight and scalable: MQTT is designed with a minimal fixed header of only 2 bytes, which reduces the communication overhead significantly. It decouples message producers and consumers, allowing easy scalability and flexibility in handling numerous sensors and devices.
  • QoS levels and reliability: MQTT’s different QoS levels enhance reliability by ensuring message delivery according to the application’s criticality. QoS 0 delivers messages at most once without acknowledgment. QoS 1 ensures messages are delivered at least once, while QoS 2 guarantees exactly one delivery through a four-step handshake, preventing duplication and ensuring data integrity for critical communications. Moreover, MQTT’s LWT (Last Will and Testament) feature allows the system to detect and respond to unexpected device disconnections, maintaining the robustness of the network.
4 Proposed DT pipeline
Figure 2: Architecture of the proposed DT pipeline: Urban Hybrid Twin.
In this section, we present a DT architecture called “Urban Hybrid Twin,” based on the sensing/communication/computing testbed deployed in NYC (Fig. 1). The proposed architecture of “Urban Hybrid Twin” is enabled by CPS, user software defined radios (SDR), cameras and LiDARs, high speed communications, and AI computing deployed in the edge cloud.
Figure 3: Diagram of intersection safety warning use case.
4.1 System components
4.1.1 Physical Infrastructure
The physical functionality of the proposed DT is built on top of the COSMOS testbed (“Cloud enhanced Open Software defined mobile wireless testbed for city-Scale deployment” [125] developed for real-world research, development, and deployment of city-scale advanced wireless communications and innovative applications [125]. It targets the technology “sweet spot” of ultra-high bandwidth and ultra-low latency, a capability that will enable a broad class of safety-time critical applications. Deployed in West Harlem, NYC, next to Columbia University campus, the COSMOS testbed is an enabler for research on the design, development, and deployment of a DT for urban traffic management.
Pilot DT experiments are executed at the intersection of Amsterdam Avenue and 120th Street, where COSMOS cameras, LiDAR, and wireless sensing and communication nodes are deployed. The cameras employ H264 encoding with an I-frame interval set to 10 frames. The live RTSP video streams–with 4K resolution at 30 fps–were processed on a COSMOS edge server equipped with an A100 GPU. The LiDAR is a Velodyne VLP-32C with up to 200m range.
Note that there is a growing trend of studies relying on emerging communication protocols like C-V2X for vehicle DTs [145, 157], which requires customized infrastructure and user-end devices. Our initial experiments utilize contemporary communications and sensing protocols for rapid deployment. They will be replaced by emerging methods that can achieve lower latency and higher bandwidths. Software defined features of the COSMOS testbed and the flexibility of radios make it possible to deploy emerging technologies such as C-V2X and experiment with novel low-latency applications.
4.1.2 Digital twin
Building on data collected from the physical world, we built an Unreal Engine based simulation platform for generating photorealistic street scenes. Moving objects are populated into an integrated SUMO-CARLA simulation platform to validate our safety warning application [43], where vehicle locations are synchronized from the real-world MQTT messages.
4.2 System functionality
4.2.1 Resource optimization
Due to the computational needs of complex computer vision models and the high volume of video data, ensuring scalability in a video analytics pipeline requires optimization of its configuration to maintain performance. However, the system’s performance is impacted by factors such as the video content, network conditions, and available computational resources. As a result, it is crucial to implement a dynamic, real-time optimization mechanism that adjusts key configuration parameters—such as resolution, frame rate, and bitrate—based on these varying conditions. This adaptive approach allows the system to continuously balance performance with resource efficiency, ensuring scalable and reliable video analytics. Accordingly, we have equipped our video analytics pipeline with a missingResource Optimization (see Fig. 2) component that continuously adapts the system’s configuration parameters to maintain its performance.
4.2.2 Use cases
4.2.2.1   Pedestrian-vehicle safety warning
Leveraging existing sensors at the intersection, we have developed a combined VRU and vehicle warning application. The cameras and the LiDAR sense the presence of VRUs at urban intersections, make predictions of their movements, apply traffic operation and control strategies, and feedback to system controllers and road users, with the primary goal of increasing traffic safety and efficiency. (see Fig. 3).
A variety of AI models have been developed to predict future trajectories of interacting road users at the intersection, including InfoSTGCAN (An Information-Maximizing Spatial-Temporal Graph Convolutional Attention Network) integrating InfoGAN and spatiotemporal graph convolutional attention network [131], and PI-NeuGODE using physics-informed graph neural ordinary differential equations[107, 110], and uncertainty quantification (UQ) to characterize the predictive confidence [105].
4.2.2.2   Learn to optimize adaptive traffic signal control
To optimize traffic signal control over a road network requires first to predict traffic patterns on road segments and intersections. The AI models we developed include PINO (physics-informed neural operator) [45] that predicts future traffic dynamics by training a mapping from initial conditions to traffic states in time and space, building on our previous work on PIDL [138, 139] and the UQ problem [108, 106]. For ATSC, we have developed an RL method, CVLight [109], and an FRL method [42] on the COSMOS network [42]. Their performances are illustrated in Fig. 4.
Figure 4: (a): ATSC for a traffic corridor. (b): Performance metrics of ATSC methods. Left: travel time and communication cost; Right: traffic volume on the incoming lane over time.
4.3 Evaluation
In safety-critical applications, it is infeasible to perform field experiments for such assessment. DT simulation is thus used to assess the effectiveness of the system.
4.3.1 End-to-end latency
Measuring latency could be particularly problematic due to the need for cross-network timing synchronization of devices and compute server. Spot measurements of several components are shown in Table 6. This table presents the latency incurred by the main components of a typical video analytics pipeline for object detection and tracking, as depicted in Fig. 3, which has been systematically optimized in terms of memory and resource usage. These elements usually run on an edge device or server. The results in Table 6 identify potential bottlenecks within the pipeline and indicate which components could benefit from further optimization. Additionally, the table illustrates the impact of video content complexity on the latency of each component, providing insights into performance variations during busy and sparse traffic conditions.
To understand the sensitivity and complexities of real-time video analytics, we measured the latency introduced by key components of a typical video analytics pipeline, illustrated in Fig. 3. For object detection, we used the YOLOv8m model [149] (batch size: 1) with the Nvidia DCF-based tracker. The YOLOv8m model was configured to accept input images at 832×832 resolution, corresponding to the size of the regions cropped by the RoI cropping component. The 5G and LTE networking latency was measured using iPhone 13 Pro Max, while wired and wireless networks provided by Columbia University Information Technology were used for the Ethernet and WiFi measurements. Latency measurements are conducted at various times of the day with different traffic densities. Tab. 6 summarizes the results of these measurements by comparing the average and standard deviation of latency per frame for the main components of the pipeline.
Metric Inference Elements Downstream Elements
Reception Pre-processing Object Detection Object Tracking MQTT Msg Creator MQTT Msg Retrieval
Ethernet Wi-Fi LTE 5G
Avg Latency (Std Dev) [ms] 1.94 (1.69) 0.108 (0.024) 7.216 (0.121) 0.973 (0.173) 0.081 (0.021) 3.21 (0.315) 6.86 (1.19) 45.72 (15.30) 39.21 (7.12)
Table 6: Average Latency and Standard Deviation for different pipeline elements.
These latency measurements demonstrate how each pipeline component influences overall latency and its variability under different video content densities. The results highlight the necessity of optimizing these components for real-time applications, particularly in environments with dynamic traffic patterns. In the following, we describe the function of each element in the inference pipeline (see Fig. 3) and their associated latencies based on the results presented in Tab. 6.
  • Reception: Receives RTP (Real-time Transport Protocol) data packets and, after parsing and decoding, converts them to raw video frames. The latency increases under busy traffic conditions, primarily due to higher complexity of the data being processed, such as larger frame sizes and more variation in the number of data packets per frame.
  • Pre-processing: Extracts regions of interest (RoIs) from raw video frames and discards irrelevant regions. It incurs low latency with minimal variations, indicating the effectiveness of pre-processing in optimizing video analytics performance.
  • Object detection: Applies the object detection model to the cropped RoIs. This element incurs the highest latency in the pipeline. The latency is slightly higher in busy traffic/crowd conditions due to more objects in each video frame.
  • Object tracking: Assigns unique IDs to detected objects across video frames for tracking. It has higher latency under busy conditions due to the increased number of objects being tracked simultaneously.
  • MQTT message creator: Converts the extracted metadata into MQTT format [104] and transmits it over a designated topic. Its latency is consistently low, with a slight increase during busy conditions due to the need to assemble larger messages.
  • MQTT message retriever: A web application captures round-trip time (RTT) of MQTT messages, and measures the latency as half an RTT.
4.3.2 Safety warning performance
Figure 5: A simulated pedestrian in CARLA (left) surrounded by cars with real-world positions (right).
To validate the accuracy of our trajectory prediction and risk assessment algorithms for warning generation, we conducted three rounds of simulation, each lasting $1010$ minutes and generating a total of $232232$ virtual pedestrians in CARLA. We then compared the number of collision warning messages with the actual number of simulated collisions. The process is illustrated in Fig. 5.
Metrics for risk assessment include time to collision (TTC) (i.e., the time remaining before a collision occurs) and post-encroachment time (PET) (i.e., the time interval between when the encroaching vehicle leaves the conflict point and when the vehicle of the right-of-way arrives at the conflict point). For prediction, we select TTC and set its threshold as $1.21.2$ seconds to minimize both false positives (FP) and false negatives (FN) while optimizing true positives (TP) and true negatives (TN). The resultant confusion matrix is $\text{TP},\text{FP}],[\text{FN},\text{TN}]]=[[66,45],[2,119]].[ [ TP , FP ] , [ FN , TN ] ] = [ [ 66 , 45 ] , [ 2 , 119$ ] ] .
5 Conclusions and open questions
Transportation applications in urban settings are generally challenging to design, develop, deploy, and test for their potential unsafe and unethical consequences. Thus, AI-empowered DT plays a critical role in effective implementation of these applications. Despite relatively sparse literature in this domain, we review an ensemble body of literature on how to leverage emerging technologies in sensing, communication, edge and cloud computing, for urban traffic management. We hope this paper can serve as a pointer to help researchers and practitioners identify challenges and opportunities on the development of DTs; a bridge to initiate conversations across interdisciplinary researchers in sensing, communication and networking, traffic engineering, and AI; a road map to exploiting potentials of DTs for diverse transportation applications.
5.1 Emerging trends
While literature on DTs for individual components has been surging, how each element of the DT works collectively and function organically is key to the development of the next-generation DT, which could empower the intelligence and automation of transportation applications.
5.1.1 Engineering the pipeline
Engineering a DT via the integration of multiple subsystems poses technical challenges, and we will name a few.
5.1.1.1   Datasets for urban computer vision
Due to limited availability and single perspective of real-world datasets, synthetic data generation from realistic 3D simulators has become increasingly popular for urban computer vision problems. The SYNTHIA dataset, for example, provides a diverse set of images from a simulated city with pixel-level semantic annotations, aiding tasks like semantic segmentation and scene understanding [129]. CARLA has been widely used in object detection and segmentation [115, 59, 101]. Similarly, GTA V has been employed to collect game frames for training autonomous agents [128]. Unreal Engine has gained traction as a rendering engine for various vision tasks [4, 55, 29]. MatrixCity utilizes Unreal Engine 5’s City Sample project to benchmark neural rendering models, focusing on pedestrian- and vehicle-free environments [78]. The Boundless project shows YOLOv8x’s enhanced performance on real-world datasets when trained with Unreal Engine 5-synthesized data [151]. However, mixed-perspective datasets that adapt to diverse deployment conditions remain scarce. Advances in open-vocabulary object detection offer a promising avenue for addressing this gap [92, 103, 24].
5.1.1.2   Sensor fusion with time synchronization
Various methods have been proposed to tackle the time-synchronization problem in multi-camera settings for a single intersection or area, either utilizing visual cues or through explicit clock synchronization between the cameras. [144] estimates the spatial transformation between the views. [9] proposes an approach at time synchronization based on the temporal alignment of matching trajectories of entities present in the overlapping scenes. [33] proposes solving a global alignment problem based on video feature descriptors. [17] uses image features to train neural networks for solving the alignment problem. [170] proposes a neural network that uses pose cues to align videos temporally. [140] proposes the use of abrupt lighting changes as temporal cues for facilitating alignment for rolling shutter cameras. Other approaches include estimations of camera capture and transmission latency [74], or clock synchronization [88, 6, 14]. These approaches address the problem of synchronizing sensors across different road intersections. Efficient, low-latency implementation of these methods are vital for synchronization and fusion of camera predictions.
5.1.2 Designing edge-cloud architecture and networking
To facilitate the development of a DT in an urban setting with numerous intersections capable of communicating with vehicles and VRUs, an extended network of cloud-connected edge devices and sensors, such as cameras, is required. These sensors generate substantial amounts of real-time data that must be transmitted to edge or cloud systems and processed with bounded latency. The data from all sensors should be integrated into a centralized platform. Given the extensive distribution of these devices, privacy and data security become critical considerations. To safeguard privacy, encoded sensor data is transmitted only to edge devices, where it is processed to extract relevant metadata. Only this metadata is then forwarded to the cloud or central platform for further analysis and integration, ensuring that no raw data or personal information is shared. To this end, data and device federation using federated (reinforcement) learning has gained growing traction [96, 35, 42].
Integrated sensing and communication (ISAC) is an emerging direction in the design of Beyond-5G wireless networks, enabling transmitted communication waveforms to be opportunistically used as radar-like sensors [89, 166]. In urban mmWave and sub-THz networks, ISAC can enable real-time tracking of vehicles and pedestrians, enhancing sensor fusion algorithms without compromising the network’s primary communication responsibilities. Future research directions may focus on leveraging high angular resolution from densely packed phased arrays to achieve precise beamforming, with the potential to introduce imaging-like capabilities within communication networks [120, 50].
5.1.3 Physics-informed DT development
The mutual interaction between domain knowledge and AI is necessary. Domain knowledge in traffic flow and road safety that has been developed for decades, provides valuable insights into every stage of the DT pipeline. In particular, it helps inform technological advancement and testbed deployment, data collection, model selection and training, resource allocation and optimization, as well as performance metrics selection. For instance, the understanding of how traffic evolves across time and space and how entities interact with one another at a road junction, could guide to what degree of granularity and precision the semantic segmentation should be done in object detection and tracking, how communication and networking resources should be prioritized, and what models and metrics should be used for risk prediction.
5.1.4 Safety guarantee of AI
DTs, heavily relying on hardware, software, and algorithms, are vulnerable to risks posed by probabilistic events. Failures in sensors, communication channels, or algorithms can result in erroneous or missing inputs to downstream prediction or decision-making, ultimately compromising the feedback loop from the DT to the physical. The rise of AI-empowered DTs introduces additional risks, including adversarial attacks, and trustworthiness in real-world applications. How can DTs be designed with provable safety guarantees, accounting for uncertainties, failures, and attacks? To achieve this, several methods have been proposed, including reachability analysis [12], robust optimization [195], and control barrier/Lyapunov functions [13, 75]. These methods are mathematically rigorous to ensure that DTs generate outputs within safe regimes. With the emergence of AI, adapting these methods to ensure the safety guarantees of AI models is becoming increasingly important. For instance, reachability analysis has been employed in safe reinforcement learning [136, 159]. Robust optimization techniques have also been applied to enhance model robustness towards adversarial attacks [19, 40]. Additionally, control barrier functions and Lyapunov functions have been integrated into neural networks to enforce safety constraints and maintain system stability [44, 22].
5.1.5 Sim-to-real gap closing
What distinguishes a traffic DT from a conventional traffic simulator? Capabilities of generating simulations and predictions consistent with the real world is the key. There is a growing amount of studies focused on sim-to-real policy learning [196, 15, 62], i.e., how optimal policies learned from a DT can be generalized to real-world implementation. The larger the sim-to-real gap, the less reliable the DT is. To the best of our knowledge, there is little research on how to close the sim-to-real gap leveraging transportation DTs, which we believe is the key to transportation applications.
5.2 Open questions: Prototyping DTs for human intelligence
Many studies define a DT as emerging technologies, namely, object detection and tracking with edge-cloud computing and communication. These technological enablers, however, are actually the “eyes” of a DT, while what really distinguishes a DT from traditional simulators lies in its “brain,” the prediction and decision making modules that are capable of extracting patterns, modeling semantics, and making informed decisions drawing upon what has been seen and perceived.
Ultimately, we should produce a next-generation of DT ideally with human-like intelligence but automation capabilities. A DT with comparable human-like intelligence or artificial artificial general intelligence (AGI) should consist of a hierarchical cognitive structure analogous to a human’s biological system, backed up by emerging hardware, software, AI algorithms, and API interfaces. The architecture of a proposed DT could consist of (see Fig. 6):
  1. 1.
    Eyes: object detection and tracking, and perception, powered by convolutional neural networks;
  2. 2.
    Brain: data storage and processing, reasoning and planning, inference and generalization, powered by AI;
  3. 3.
    Neural systems: edge-cloud networking and computing backbone, powered by resource allocators;
  4. 4.
    Communication: natural language processing, powered by large language models (LLM) and foundation models.
Figure 6: Schematic diagram of a DT analogous to humans.
A growing number of sensors and explosive amounts of data have opened up tremendous opportunities for researchers to apply AI to create, train, evaluate, and streamline traffic DT pipelines, from object detection, tracking, prediction, simulation, to traffic operation and management. The emerging LLM and foundational models have further demonstrated great potential in DTs. LLMs could generate new data for training data augmentation or privacy concerns [165, 191, 21], enhance interactions between human users and the DT system [188], making personalized recommendations [177, 38] and even automate the creation process of DTs [10].
Despite the promising future of AI-powered DTs, application of AI to DTs could face challenges. As opposed to classical statistical methods, AI algorithms are generally black boxes where their inner workings are not transparent to developers nor users. Thus, interpretability or explainability using Shapley value [100], PIDL [32], symbolic regression [102], or Kolmogorov-Arnold networks [97] can potentially reveal to some degree the rationale underlying the predictions. Unlike humans, AI models do not understand causes and heavily rely on correlations to make predictions. Without knowing causality could render AI methods incapable of generalizing to unseen data. Thus, augmenting AIs with causal reasoning and inference could increase its deductive capabilities [130, 132, 133]. In addition, generative AIs could produce hallucination, which might lead to nonphysical predictions or unrealistic decision making. Introducing physics based domain knowledge could help fine tune these models, enhance inductive biases, and generate more meaningful outputs.
Last but not the least, there has been a growing trend in developing safe AI systems aligned with human values and objectives. For example, recent studies have focused on evaluating LLMs in terms of toxicity [25], privacy [180], ethics [119], and fairness [111], indicating that LLMs are not sufficiently safe. It is thus crucial to continually achieve safety guarantees of AI, especially for the implementation of DTs.
Acknowledgment
We would like to thank Mengxuan Liu for her assistance in generating the 3D models.

### v2

@@
-AI-Powered Urban Transportation Digital Twin: Methods and Applications
+AI-Powered CPS-Enabled Urban Transportation Digital Twin: Methods and Applications
 Abstract
-We present a survey paper on methods and applications of digital twins (DT) for urban traffic management.
+We present methods and applications for the development of digital twins (DT) for urban traffic management.
 While the majority of studies on the DT focus on its “eyes,” which is the emerging sensing and perception like object detection and tracking, what really distinguishes the DT from a traditional simulator lies in its “brain,” the prediction and decision making capabilities of extracting patterns and making informed decisions from what has been seen and perceived.
-In order to add values to urban transportation management, DTs need to be powered by artificial intelligence and complement with low-latency high-bandwidth sensing and networking technologies.
-We will first review the DT pipeline leveraging cyberphysical systems and propose our DT architecture deployed on a real-world testbed in New York City.
-This survey paper can be a pointer to help researchers and practitioners identify challenges and opportunities for the development of DTs; a bridge to initiate conversations across disciplines; and a road map to exploiting potentials of DTs for diverse urban transportation applications.
-footnotetext: Xuan Di, Yongjie Fu, Mehmet K.Turkcan, and Zhaobin Mo are with the Department of Civil Engineering and Engineering Mechanics, Columbia University.
-Emails: {sharon.di, yf2578, mkt2126, zm2302}@columbia.edufootnotetext: Xuan Di, Zoran Kostic, and Gil Zussman are also with the Data Science Institute, Columbia University.footnotetext: Mahshid Ghasemi, Chengbo Zang, Abhishek Adhikari, Zoran Kostic, and Gil Zussman are with the Department of Electrical Engineering, Columbia University.Emails: {mg4089, cz2678, abhishek.adhikari, zk2172, gil.zussman}@columbia.edu
-Keywords: Digital twin, AI, Urban transportation management, Survey
-1 Introduction
+In order to add value to urban transportation management, DTs need to be powered by artificial intelligence and complement with low-latency high-bandwidth sensing and networking technologies, in other words, cyberphysical systems (CPS).
+We will first review the DT pipeline enabled by CPS and propose our DT architecture deployed on a real-world testbed in New York City.
+This paper can be a pointer to help researchers and practitioners identify challenges and opportunities for the development of DTs; a bridge to initiate conversations across disciplines; and a road map to exploiting potentials of DTs for diverse urban transportation applications.
+Index Terms:
+Digital twin, AI, Urban traffic management
+I Introduction
 Urban transportation systems are complex to model and simulate, due to heterogeneous road users (such as cars, pedestrians, cyclists, scooters) interacting in multimodal traffic environments consisting of public and private travel modes.
-With fast-changing traffic evolution in time and space, traffic simulation if improperly calibrated, might produce traffic management strategies that largely deviate from the reality, potentially leading to suboptimal or even detrimental outcomes.
+With fast-changing traffic evolution in time and space, traffic simulation, if improperly calibrated, might produce traffic management strategies that largely deviate from the reality, potentially leading to suboptimal or even detrimental outcomes.
 With ubiquitous sensors in smart cities, it is the time to augment conventional traffic simulators, many of which were developed in an era when only “small data” became available.
 Emerging traffic sensors are expected to generate big volumes of data, transmitted via communication networks and processed on edge cloud computing with artificial intelligence (AI) for real-time traffic management.
-Such a transformation poses challenges that call for the development of a new paradigm, namely, digital twin (DT), which will push the envelope in urban transportation management.
+Such a transformation calls for the development of a new paradigm, namely, digital twin (DT), which will push the envelope in urban transportation management.
 Literally, DT is the digital replica of a physical object or asset [1], where a digital world mirrors a physical world for real-time diagnosis, prognosis, and decision making.
-Recent years have seen a growing amount of studies on DT in various domains, ranging from manufacturing [48], aerospace [76], to smart cities [82].
-In particular, there is a sizable body of articles on transportation DTs powered by vehicular technologies such as connectivity and autonomy [134, 162, 57, 199, 80].
-With such a growing trend, we would like to restrict the scope of this survey paper.
-Here we are particularly interested in urban traffic settings where cars have to interact with vulnerable road users (VRU) (i.e., non-motorists such as pedestrian, bicyclists, other cyclists, or persons on personal conveyance [153]), who are an integral part and also potential users of the DT.
-We will primarily focus on two use cases, namely, intersection safety warning and network wide traffic signal control, for improved traffic safety and efficiency.
-The literature on DTs for urban traffic management is lacking, partly because the development of a DT for a system is non-trivial, particularly involved with human participants like VRUs.
-The development of such a DT must engage with expertise in sensing, communication, computing, and human-centric perspectives.
-Accordingly, we strongly believe that a comprehensive review is warranted on trans-disciplinary topics.
-This paper presents a survey paper on methods and applications of DT for urban traffic management.
+Recent years have seen a growing amount of studies on DT in various domains [1, 2, 3], including a sizable body of articles on vehicular DTs [4, 5, 6, 7, 8].
+With recent explosive growth of literature on DTs, we would like to restrict the scope of this paper to applications in the urban setting, especially when vulnerable road users (VRU) (i.e., non-motorists such as pedestrian, bicyclists, other cyclists, or persons on personal conveyance [9]) are an integral part of the system and also potential users of the DT.
+We will primarily focus on use cases accounting for VRUs with improved traffic safety and efficiency.
+The studies of DTs for urban traffic management, especially involving VRUs, are lacking, partly because the development of a DT for a system is non-trivial, particularly when involved with humans.
+This paper presents methods and applications for the development of DTs for urban transportation systems.
 We will depict a DT pipeline prototype, leveraging the architecture of cyberphysical systems and AI methods.
-We will propose a DT for real-time traffic monitoring and optimization [127], based on an existing physical testbed deployed in New York City (NYC) leveraging cutting-edge sensing, communication, computing, and AI-based automation.
+We will propose a DT for real-time traffic monitoring and optimization, based on an existing physical testbed deployed in New York City (NYC) leveraging cutting-edge sensing, communication, computing, and AI-based automation.
 The overall contributions of this paper include: (1) introducing AI methods used in the DT pipeline; (2) exploring the architecture of transportation DTs and propose a prototype for reference; and (3) identifying gaps and directions.
 The rest of this paper is organized below.
-Section 2 unifies various concepts of DT and introduces the relationship between cyberphysical transportation systems and DTs; Section 3 reviews the literature on sensing and communication technologies, object detection and tracking, and urban transportation use cases.
-Section 4 demonstrates the architecture of our proposed DT building on a real-world testbed.
-Section 5 concludes our work and presents potential research directions as well as open questions.
-2 Preliminaries
-2.1 Variants of DT definitions
+Section II introduces transportation DTs enabled by cyberphysical systems, and position this paper; Section III reviews the literature along the pipeline of a DT.
+Section IV demonstrates the architecture of our DT, building on a real-world testbed.
+Section V concludes our work, presents potential research directions and open questions.
+II Preliminaries
 The definition of DT has evolved rapidly.
-Despite presenting its own version, these articles all share common elements and resemble certain characteristics [48, 114].
+Despite presenting its own version, articles share common elements and resemble certain characteristics [1, 10, 11].
 In general, there is a physical world (aka.
@@
 In other words, the ultimate goal of a DT development is to add values to the physical for improved safety and efficiency.
-Depending on whether there is any interaction between the physical and the digital via data and/or information exchange, there are three types of digital twinning concepts [47], namely,
-  1. 1.
-    digital model: with no automatic data exchange;
-  2. 2.
-    digital shadow: with only one-way data exchange from the physical to the digital;
-  3. 3.
-    digital twin: with two-way communication.
-Let us specify a definition of DT in transportation.
-Definition 2.1
-Traffic digital twin is a digital system integrating the pipeline from object detection and tracking, resource allocation, edge-cloud computing and communication, for online simulation, operation, control and management.
+Below, we first formalize transportation DT, and then discuss the relation between DT and cyberphyscial systems.
+Definition II.1
+Transportation digital twin (T-DT) is a digital system integrating the pipeline from object detection and tracking, resource allocation, edge-cloud computing and communication, for online simulation, operation, control and management.
 It is updated online using continuously fed data collected from the physical and send control policies or issue warnings back to the physical, leveraging big data and AI tools.
-Traffic DTs are closed-loop with two-way communication, where data, information, and control signals are exchanged with the physical sequentially and reiteratively.
-2.2 Cyberphysical transportation systems
-Definition 2.2
+T-DTs are closed-loop with two-way communication, where data, information, and control signals are exchanged with the physical sequentially and reiteratively.
+Definition II.2
 Cyberphyscial systems (CPS) [12] are smart systems that include engineered interacting networks of physical and computational components.
 CPS holds great potential to enable real-time applications thanks to emerging technologies in sensing, communication, and computing.
-2.2.1 CPS as DT technological enabler
 A transportation CPS interlinks physical and cyber layers, where the cyber layer consists of sensing, networking, computing, and traffic management application modules (see Fig.
@@
 To enable the technological development of a DT, a physical testbed is needed as a platform for sensing, computing, experimentation, evaluation, as well as design constraints determination.
-In Secs.
-3.1-3.3, we will review the CPS technological enablers needed for the development of a DT, and examine the testbed used for our proposed DT in Sec.
-4.1.1.
-2.2.2 DT as CPS testing tool
-An automatic system like CPS demands high degree of autonomy for urban traffic management, with increased complexity and scale.
-Thus, DT becomes crucial for the test and verification of a CPS.
-We need to verify the system behavior of a CPS to ensure that it performs as desired [141], such as safety, efficiency, timeliness, and accuracy.
-Due to the integration of cyber and physical components, as well as their two-way coupling via communication and networking, testing a CPS is non-trivial.
-Given that the digital is expected to be the twin of the physical, it is natural to compare the difference in the outputs of both the physical and the DT using “Grieves performance test” [48].
-Hardware-in-the-loop testing, such as vehicle-in-the-loop [172], could help test directly the performance of some physical components “live,” while allowing the rest to be simulated within a DT.
-For DTs that involve the feedback to humans, human-in-the-loop test is widely used on driving simulators [183, 79].
-Recent years have seen a growing trend of using augmented reality and virtual reality (AR/VR) or Metaverse to engage humans as pedestrians, cyclists, or scooters in virtual environments without inducing real-world risks [164, 72].
-Since there does not exist a unified approach for the DT testing in the context of transportation, in Sec. 3.4, we will review evaluation metrics employed by various use cases.
-3 Related work
-In this section, we will review technological enablers of DT along the pipeline depicted in Fig. 1.
-Figure 1: Pipeline of a CPS-structured DT.
-3.1 Sensing and perception
+In Sec.
+III, we will review the CPS technological enablers needed for the development of a DT, and examine the testbed used for our proposed DT in Sec.
+IV.
+We would like to stress that, this paper aims to discuss how AI algorithms and the architecture of CPS contribute to the urban T-DT (UT-DT) pipeline.
+In contrast, there are highly cited survey papers on T-DT, which are more focused on general transportation scenarios and applications, while AI might not be the key focus.
+Tab.
+I outlines the comparison of a partial set of related work.
+TABLE I: Comparison of survey papers on DT
+Ref. Topic Focus
+[4][5][6] Comprehensive review for vehicle mobility apps Vehicular technology driven pipeline
+[13] Comprehensive review for traffic safety and mobility CPS pipeline highlighting communication
+[14] Comprehensive review for operation and maintenance apps Broad travel modes
+Ours Semi-technical review and position for urban traffic apps AI-Powered CPS pipeline
+We first review various CPS and AI methods needed in the pipeline, and then present our T-DT instance.
+Accordingly, this paper is semi-survey, semi-technical.
+We employ such an organization, because we find that the publications on T-DT are normally segmented by different communities and journals, which could prevent researchers from understanding the entire pipeline, from upstream sensing and perception, to downstream transportation applications.
+For example, [15] on networking DT is primarily focused on the development of communication technology with evaluation on latency, even its use case is transportation.
+Transportation researchers, who hope to implement this system in real-world, have to seek more details about how traffic dynamics would impact the performance, which is unfortunately missing.
+This is likely because the authors belong to the society of communication and networking.
+Another example is that, [3] focused on vehicular DT, heavily rely on the foundational knowledge in CPS, which could be somewhat unfamiliar to many transportation researchers.
+Realizing such a gap in a large body of literature on T-DT, this paper aims to unify the knowledge by presenting a comprehensive summary upfront, and exemplify the pipeline following the summary.
+After all, the development of a T-DT calls for the interdisciplinary collaboration across transportation, electrical engineering, mechanical engineering, computer science, and human-machine interaction.
+To foster the readability of the paper, in the next section, we first offer a comprehensive state-of-the-art review on related work along the CPS pipeline.
+III Related work on CPS-enabled DT
+The development of CPS-enabled DT must engage with expertise in sensing, communication, computing, and human-centric perspectives, where AI methods are backbones.
+Figure 1: Schematic diagram of a DT.
+III-A Sensing and perception
 Sensors are the “eyes” (and “ears”) of a DT.
 Tab.
-2 summarizes the pros and cons of each sensing technology, namely, mobile devices, on-board vehicles, and roadside infrastructure, for urban traffic applications.
+II summarizes the pros and cons of each sensing technology, namely, mobile devices, on-board vehicles, and roadside infrastructure, for urban traffic applications.
 The former two are mobile sensors with wider spatial coverage but challenge in precision because of moving references, while the latter at fixed locations could face limited sensing ranges and coverage.
-Table 1: Pros and cons of sensors for urban traffic applications (partially adapted from [154, 194])
+TABLE II: Sensors for urban traffic applications (partially adapted from [16, 17])
 Sensor Purpose Advantages Disadvantages
@@
 3. Limited field of view when the range is far [19].
-Camera Obstacle and lane detection 1. Maintains good resolution when the FOV is wide.
+Camera Obstacle and lane detection 1. Maintains good resolution when the field of view is wide.
 2. Has a long horizon.
@@
 Accelero -meter Acceleration, driving behavior 1. Detects braking, turning, and accelerating accurately.
-2. Can be integrated with the vehicle system for behavior analysis [98].
+2. Integrated with vehicles for behavior analysis [20].
 1. Performs poorly for slow or subtle movements.
@@
 3. Has difficulty with real-time calibration [23, 24].
-3.2 Object detection and tracking
+III-B Object detection and tracking
 Object detection identifies and classifies objects within the environment using sensors like cameras, Lidar, and Radar.
@@
 Here we thus summarize commonly used and emerging datasets from non-vehicle sensors in Tab.
-1.
-Table 2: Public datasets for urban traffic scenes (On-board vehicle datasets can be found in [31].)
+III.
+TABLE III: Public datasets for urban traffic scenes (On-board vehicle datasets can be found in [25].)
 Sensor location Dataset Purpose Sensor Setup Collection region
@@
 DeepSense 6G [29] 2D/3D object detection, sensor fusion RGB cameras, mmWave Phase Arrays, LiDAR, Radar Various indoor and outdoor spaces
-Aerial VisDrone [202], NGSim [116], highD [67], rounD [68], DOTA [171], CitySim [198] Aerial object detection and tracking, trajectory forecasting Drone-based RGB cameras Urban spaces in China and Aacheny, intersections in California and Florida
+WILDTRACK [30] Multi-Object Tracking RGB cameras University campus in Zurich
+Aerial VisDrone [31], NGSim [32], highD [33], rounD [34], DOTA [35], CitySim [36] Aerial object detection and tracking, trajectory forecasting Drone-based RGB cameras Urban spaces in China and Aachen, intersections in California and Florida
+MOT Challenge [37], MOTS20 [38], UAVDT [39] Aerial object detection and tracking, trajectory forecasting Drone-based RGB cameras Various Indoor and Outdoor Scenes
 Misc CoCo [40], ADE20K [41], Cityscapes [41] Object detection, semantic segmentation RGB cameras Various indoor and outdoor spaces
+Synthetic CARLA [42, 43, 44] Autonomous-driving object detection and segmentation RGB cameras, LiDAR Urban European/North American environments
+GTAV [45], Synscapes [46], UrbanSyn [47] Object detection and segmentation RGB cameras, LiDAR Urban European/North American environments
+MOTSynth [48] Multi-object tracking RGB cameras Urban European/North American environments
+MatrixCity [49] Neural-rendering benchmarking (vehicle/pedestrian-free) RGB cameras Synthetic city environment
+Boundless [50] Object detection and segmentation with UE5-synthesised data RGB cameras Synthetic urban environments
 Object detection has been studied extensively for urban applications.
+A large number of studies focus on low-altitude vehicle and pedestrian detection [51, 52].
+Many focus on high-altitude aerial environments, where small object detection becomes an important challenge [31, 35, 33, 36].
 Single-stage object detectors, following the original single shot multibox detector [53] and You Look Only Once (YOLO) [54] architectures, have become popular due to their real-time deployment capabilities.
-In the last few years, transformer-based object detection approaches, competitive with YOLO models, have emerged as the state-of-the-art in object detection when designed to be deployed in real-time settings [197].
+In the last few years, transformer-based object detection approaches, competitive with YOLO models, have emerged as the state-of-the-art in object detection when designed to be deployed in the real-time setting [55, 56, 57, 58].
 Recent progress in YOLO object detection performance has been enabled through multiple small tricks in architecture and training that all together provide significant improvements in empirical performance [59].
@@
 BoostTrack explores novel distance and shape similarity measurements to deal with ambiguity caused by unreliable detection results [76, 77].
-Modern MOT pipelines are often implemented directly downstream to object detection in the tracking-by-detection paradigm.
+Predicting future trajectories of detected objects is often a crucial part of safety-critical applications.
+Numerous deep neural network models have emerged as competitive candidates for trajectory prediction over the past few years.
 Majority of modern architectures for predicting future trajectories of detected objects adopt Recurrent Neural Networks which is responsible for predicting future object positions based on their historical coordinates, together with generative components which handles the variation and flexibility in social interactions [78, 79].
@@
 Unlike the tracking algorithms, specific training or fine-tuning is often required before the deployment of trajectory forecasting models to unseen scenarios.
-3.3 Real-time video analytics
+III-C Real-time video analytics
 Developing end-to-end real-time video analytics systems on a large scale for time-sensitive and safety-critical traffic and crowd management applications presents challenges.
@@
 Tab.
-3 provides a comparative analysis of various approaches to address these challenges.
+IV provides a comparative analysis of various approaches to address these challenges.
 The approaches differ in their focus–some prioritize reducing latency and resource consumption, while others emphasize maintaining or enhancing accuracy, especially under constrained conditions.
 This comparison provided in Tab.
-3 highlights the trade-offs inherent in real-time video analytics and emphasizes different optimization methods to balance throughput, accuracy, energy consumption, and computational efficiency across diverse deployment scenarios, including edge devices, cloud platforms, and hybrid environments.
-Table 3: Comparison of various approaches and optimization objectives in video analytics.
+IV highlights the trade-offs inherent in real-time video analytics and emphasizes different optimization methods to balance throughput, accuracy, energy consumption, and computational efficiency across diverse deployment scenarios, including edge devices, cloud platforms, and hybrid environments.
+TABLE IV: Comparison of various approaches and optimization objectives in video analytics.
 References Approach Optimization objective
@@
 Minimize energy consumption for CNN inference services on high-performance GPUs while meeting latency of Service-Level Objectives.
-3.4 Use cases
+III-D Communication and networking
+A DT for safety-critical applications requires real time communications with aggressively low latency.
+We explore issues related to low latency communications, and survey component technologies and protocols that can be utilized to achieve very low latency.
+III-D1 Real-time requirements and low latency targets
+Sensor and control data in a real time system is subject to latency created by the stages, namely, (1) data acquisition from traffic participants (such as camera recordings and encoding, harvesting data from autonomous vehicles, and collect information from fiber); (2) transmission of data across communications links from sensors to inferencing servers using communications protocols such as Transmission Control Protocol (TCP), Unreliable Data Protocol (UDP) and Real-Time Streaming Protocol (RTSP); (3) data preprocessing (video decoding, and cropping); (4) AI inferencing; (5) higher level reasoning about required feedback to traffic participants; and (6) sending feedback to traffic participants across communications links via low-latency broadcast or dedicated channels over wired and wireless.
+Smart city applications can be grouped according to their latency requirements.
+Many, if not all, pedestrian-associated application (facilitated by pedestrian detection/observations and message notifications) are likely to expect the round trip delay in the range of a couple of seconds.
+Such latency can be supported by contemporary cameras, communication protocols and inferencing engines.
+Applications which would attempt to close the observation/notification loop for vehicles moving at about 10 km/h may expect latency in tens of millisecond.
+Using conventional video compression, RTP/RTSP streaming and edge computing is inadequate to support such latency.
+This presents the opportunity to pursue novel engineering solutions and research problems.
+III-D2 Communication techniques and protocols
+Ultra-Reliable Low Latency Communications (URLLC), a key component of 5G wireless, can help achieve the low latency targets.
+Along with enhanced mobile broadband (eMBB) and massive machine-type communication (mMTC), URLLC [115] represents one of the three main capabilities of 5G New Radio (5G NR), as standardized by the 3rd Generation Partnership Project (3GPP).
+In the context of transportation systems, URLLC aims to deliver up to 99.999% reliability and single-digit millisecond latency [116].
+However, meeting these performance metrics is challenging in practice due to complex channel environments, particularly in dense urban areas, which can reduce reliability.
+For intelligent transportation systems, where URLLC may be used as infrastructure backhaul, the target is an end-to-end latency of 30 ms [117].
+An emerging technology in this space is Cellular-Vehicle-to-Everything (C-V2X), which has largely replaced the earlier Wi-Fi-based Dedicated Short-Range Communications (DSRC).
+Unlike DSRC, C-V2X leverages cellular networks, allowing network providers to offer always-on connectivity, which is a critical feature for time-sensitive applications.
+Additionally, private 5G networks are being developed to ensure this level of connectivity, overcoming the congestion and range limitations inherent to Wi-Fi [118].
+In transportation systems, active collaboration between wireless service providers and vehicle manufacturers is in progress to integrate private 5G networks into vehicular networks [119].
+Tab.
+V summarizes the key characteristics of commonly used IoT communication protocols, Message Queuing Telemetry Transport (MQTT) [120], Constrained Application Protocol (CoAP) [121], and Hypertext Transfer Protocol (HTTP) [122].
+Among them, MQTT emerges as a practical choice, due to its combination of low latency, high scalability, and reliable delivery mechanisms.
+Its lightweight publish/subscribe model, Quality of Service (QoS) guarantees, and session management make it well suited for real-time data exchange in dynamic, safety-critical environments like urban transportation systems [123, 124].
+TABLE V: Communication protocol comparison for real-time digital twin systems.
+Protocol Latency Scalability Reliability Best Suited For
+MQTT Persistent TCP connection delivers low latency for ongoing messaging Highly scalable with a broker-based publish/subscribe model supporting many-to-many communication across thousands of devices Guarantees reliable delivery with configurable acknowledgment levels, exactly-once delivery via a four-step handshake, persistent sessions Large-scale IoT, real-time sensor/actuator networks
+CoAP UDP-based, very low latency when network is stable, but performance degrades with packet loss Limited scalability due to client-server request/response model.
+Multicast is possible but complex and unreliable at scale Basic two-way acknowledgment, no exactly-once delivery guarantee, no session continuity Resource-limited devices that send small, infrequent data
+HTTP Higher latency due to new TCP/TLS handshake and headers per request Limited scalability due to client-server request/response model and verbose ASCII headers Relies on TCP for delivery, with no application-level acknowledgment, retries, or session handling Web APIs, periodic data transfer, backend integration
+IV Proposed DT pipeline
+Figure 2: Architecture of the proposed DT pipeline: Urban Transportation DT (UT-DT).
+In this section, we present a DT architecture for UT-DT, based on the sensing/communication/computing testbed deployed in NYC (Fig.
+2).
+The proposed architecture is enabled by cameras and LiDARs, high speed communications, and edge cloud computing.
+Figure 3: Diagram of intersection safety warning use case.
+IV-A Physical Infrastructure
+Pilot experiments are executed at the signalized intersection of Amsterdam Avenue (major) and 120th Street (minor) near Columbia campus in NYC.
+The road geometry and traffic statistics are summarized in Fig.
+4.
+Figure 4: Traffic statistics for NYC’s intersection of 120th St. & Amsterdam Ave.
+IV-A1 Technological enabler
+The physical functionality of the proposed UT-DT is built on top of the COSMOS testbed (“Cloud enhanced Open Software defined mobile wireless testbed for city-Scale deployment”) [125], developed for real-world research, development, and deployment of city-scale advanced wireless communications and innovative applications [125].
+It targets the technology “sweet spot” of ultra-high bandwidth and ultra-low latency, a capability that will enable a broad class of safety-time critical applications.
+Deployed in West Harlem, NYC, next to Columbia University campus, the COSMOS testbed is an enabler for research on the design, development, and deployment of a DT for urban traffic management.
+At the intersection of Amsterdam Avenue and 120th Street, cameras, LiDAR, and wireless sensing and communication nodes are deployed.
+The cameras employ H264 encoding with an I-frame interval set to 10 frames.
+The live RTSP video streams–with 4K resolution at 30 fps–were processed on a COSMOS edge server equipped with an A100 GPU.
+The LiDAR is a Velodyne VLP-32C with up to 200m range.
+IV-B Digital twin
+Building on data collected from the physical world, we built an Unreal Engine based simulation platform for generating photorealistic street scenes.
+Moving objects are populated into an integrated SUMO-CARLA simulation platform to validate our safety warning application [126], where vehicle locations are synchronized from the real-world MQTT messages.
+[…]
+(a) COSMOS in CARLA
+(b) COSMOS in SUMO
+(c) Boundless with snow
+(d) Boundless with rain
+Figure 5: (Top) Carla-SUMO simulator; (Bottom) Qualitative examples from Boundless. Bounding boxes for vehicles are shown in blue, and pedestrians are shown in green.
+IV-C Use cases
 The value of a DT for urban transportation management ranges from traffic state prediction, spatiotemporal traffic flow forecasting, to urban planning and policy making.
-Here we focus on two that are particularly important in urban settings.
-3.4.1 VRU safety warning: Safety-critical application
+Here we focus on two that are particularly important in urban settings, leveraging video analytics.
+IV-C1 Use case 1: Micromobility detection
+In the U.S., the e-bike had a market size of almost \$2 billion in 2022, with a projected growth of 15.6% from 2023 to 2030.
+In 2022, 1.1 million e-bikes were sold, four times as many as were sold in 2019.
+In 2019, 136 million trips were made using micromobility, a 60% increase from 2018.
+Because their presence and surge have profound implications for road safety and urban traffic management, automatic detection of micromobility for regulation becomes increasingly important.
+However, there does not exist a standard AI model nor a dataset for the object detection of this emerging transportation mode.
+AI model evaluation
+[…]
+TABLE VI: Class-wise recall (%) on the Micromobility test set. Higher values indicate better detection performance.
+Model Bicycle $\uparrow$ e-Bike $\uparrow$ Motorcycle $\uparrow$
+YOLOv9e 52.9 54.7
+Ours 68.1 61.0 91.8
+Figure 6: Four sample frames showing the small scale of micromobility instances from infrastructure-based 4K resolution cameras in the COSMOS testbed.
+IV-C2 Use case 2: VRU safety warning
 Safety-critical applications, making automated life-and-death decisions such as collision avoidance warning between automobiles and pedestrians, need to activate at the precise time and the right moment with bounded latency.
@@
 The applications of DTs on safety-critical scenarios, however, remains understudied, because it necessities short runtime and real-time reaction, posing high requirements for communication and computing technologies, pipeline architecture design, and testing.
-We summarize the safety-critical applications to address conflict risks between vehicles and VRUs in Tab. 4, and offer an outlook of safety guarantee related methods in Sec. 5.1.4.
-Table 4: Literature review on safety warning to VRUs.
+We summarize the safety-critical applications to address conflict risks between vehicles and VRUs in Tab.
+VII, and offer an outlook of safety guarantee related methods in Sec.
+V-A3.
+TABLE VII: Literature review on safety warning to VRUs.
 Ref Technology Objective Risk Assessment Metrics Evaluation Method/Metric
@@
 Computing: Local server equipped with GPU
-Develops a video-based vehicular BSMs method with lower error and latency that outperforms the cellular vehicle-to-everything (C-V2X) method. N/A 1. End-to-end delay: <100absent100<100< 100  ms
+Develops a video-based vehicular BSMs method with lower error and latency that outperforms the cellular vehicle-to-everything (C-V2X) method. N/A 1. End-to-end delay: <100<100  ms
 2. Localization/Speed accuracy
@@
 Computing: CAM server deployed at edge/cloud
-Proposes a system using commercial devices and standard messages for road user communication. The distance between VRUs and nearby entities Latency from VRU to CAM <50absent50<50< 50 ms
+Proposes a system using commercial devices and standard messages for road user communication. The distance between VRUs and nearby entities Latency from VRU to CAM <50<50 ms
 [131] Sensing: Camera, Android phone’s GPS module
@@
 Computing: Edge computing server, smartphones
-[…]
+Develops a system to deliver CAM to VRUs on smartphones using Beacon stuffing without the need for root access to utilize 802.11p. N/A 1. End-to-end latency $\sim$ 2500 ms
 2. MAC channel utilization
@@
 The distance between VRUs and nearby entities 1.
-End-to-end latency <300absent300<300< 300 ms
+End-to-end latency <300<300 ms
 2. Distance error of vehicles and VRUs
@@
 2. Granular latency per frame
-3.4.2 Networkwide adaptive traffic signal control
-Learning methods for network wide adaptive traffic signal control (ATSC), including reinforcement learning (RL) and federated learning (FL), aim to learn optimal traffic signal control strategies in traffic environments with unknown state transitions and unknown rewards.
-Assuming that traffic signal controllers are intelligent autonomous agents, their goal is to make sequential decisions to optimize an objective function, such as minimization of traffic delay or maximization of traffic throughput.
-To train an optimal policy for traffic signal controllers, a traffic simulator or DT, needs to be built, where dynamic traffic is simulated in the training environment aligned with the testing environment in principle.
-The trained policies leveraging such a DT need to be executed in the real world, which triggers updated environments for future decision making.
-Related work are summarized in Tab.
-5.
-Table 5: Summary of papers using (F)RL for ATSC
-Ref. RL FL DT Simulator Real-world Demand Real-world Network
-[63] DDPG SUMO
-[30] SUMO
-[71] A2C Vissim
-[64] Vissim
-[201] SUMO
-[42] A3C FedAvg SUMO
-[182] DQN FedAvg Cityflow
-[58] PPO FedAvg SUMO
-[16] DQN FedAvg SUMO
-[109] A2C SUMO
-[77] A2C SUMO
-[27] A2C SUMO
-[52] PPO SUMO
-FedAvg: Federated Averaging, DDPG: Deep Deterministic Policy Gradient, A2C: Advantage Actor Critic, A3C: Asynchronous Advantage Actor Critic, PPO: Proximal Policy Optimization, DQN: Deep Q Learning.
-3.5 Close the loop
-To realize the value of a DT, the feedback from the digital to the physical is the key. The type of feedback and to what entities would define the technological constraints.
-  • Feedback to VRUs (super local): When users react to an issued warning, the state of the environment for the VRU would change, and the DT could update its status about whether continuous warning is needed.
-Normally the warning is safety related, thus demanding low latency and high bandwidth.
-  • Feedback to cars (super local): The feedback to cars includes hierarchies: (1) strategic driving feedback, such as routing recommendations, allowing higher latency and lower bandwidth; (2) tactical feedback on tactical driving decisions, such as lane-change, with medium latency and bandwidth; and (3) operational feedback such as throttle control, demanding (ultra) low latency and bandwidth.
-When cars change their driving maneuvers, the traffic environment needs to be updated and unsafe outcomes could be potentially avoided.
-  • Feedback to traffic operators and infrastructure (local or networkwide): For traffic operators, DTs could provide anomaly detection for diagnosis, and predict future traffic states or collisions for prognosis, which would help with variable speed limit, dynamic speed advisory, lane usage adaptation, ramp metering, traffic signal timing optimization, and network wide traffic management, with increasing requirements for latency and bandwidth.
-  • Feedback to city planners (citywide): In the long-term, a DT could provide aggregate traffic information to city planners to assess the potential societal impact of new infrastructure planning.
-Thus, the technological requirement is relatively low, but the collected data needs to cover a relatively large spatiotemporal range for citywide decision making,
-3.6 Communication and networking
-A DT for safety-critical applications requires real time communications with aggressively low latency.
-We explore issues related to low latency communications, and survey component technologies and protocols that can be utilized to achieve very low latency.
-3.6.1 Real-time requirements and low latency targets
-Sensor and control data in a real time system is subject to latency created by the stages, namely, (1) data acquisition from traffic participants (such as camera recordings and encoding, harvesting data from autonomous vehicles, and collect information from fiber); (2) transmission of data across communications links from sensors to inferencing servers using communications protocols such as Transmission Control Protocol (TCP), Unreliable Data Protocol (UDP) and Real-Time Streaming Protocol (RTSP); (3) data preprocessing (video decoding, and cropping); (4) AI inferencing; (5) higher level reasoning about required feedback to traffic participants; and (6) sending feedback to traffic participants across communications links via low-latency broadcast or dedicated channels over wired and wireless.
-Smart city applications can be grouped according to their latency requirements.
-Many, if not all, pedestrian-associated application (facilitated by pedestrian detection/observations and message notifications) are likely to expect the round trip delay in the range of a couple of seconds.
-Such latency can be supported by contemporary cameras, communication protocols and inferencing engines.
-Applications which would attempt to close the observation/notification loop for vehicles moving at about 10kmph may expect latency in tens of millisecond.
-Using conventional video compression, RTP/RTSP streaming and edge computing is inadequate to support such latency.
-This presents the opportunity to pursue novel engineering solutions and research problems.
-3.6.2 Communication techniques and protocols
-Ultra-Reliable Low Latency Communications (URLLC), a key component of 5G wireless, can help achieve the low latency targets.
-Along with enhanced mobile broadband (eMBB) and massive machine-type communication (mMTC), URLLC [123] represents one of the three main capabilities of 5G New Radio (5G NR), as standardized by the 3rd Generation Partnership Project (3GPP).
-In the context of transportation systems, URLLC aims to deliver up to 99.999% reliability and single-digit millisecond latency [2].
-However, meeting these performance metrics is challenging in practice due to complex channel environments, particularly in dense urban areas, which can reduce reliability.
-For intelligent transportation systems, where URLLC may be used as infrastructure backhaul, the target is an end-to-end latency of 30 ms [1].
-An emerging technology in this space is Cellular-Vehicle-to-Everything (C-V2X), which has largely replaced the earlier Wi-Fi-based Dedicated Short-Range Communications (DSRC).
-Unlike DSRC, C-V2X leverages cellular networks, allowing network providers to offer always-on connectivity, which is a critical feature for time-sensitive applications.
-Additionally, private 5G networks are being developed to ensure this level of connectivity, overcoming the congestion and range limitations inherent to Wi-Fi [37].
-In transportation systems, active collaboration between wireless service providers and vehicle manufacturers is in progress to integrate private 5G networks into vehicular networks [3].
-Message Queuing Telemetry Transport (MQTT) is a lightweight communication protocol that offers efficient, reliable, and scalable data transfer necessary for real-time, safety-critical applications.
-Such a protocol can efficiently collects and distributes traffic data from numerous sensors to update DTs in real-time, and facilitates fast communication of incidents and emergencies, allowing timely interventions.
-Key features include [122]:
-  • Lightweight and scalable: MQTT is designed with a minimal fixed header of only 2 bytes, which reduces the communication overhead significantly.
-It decouples message producers and consumers, allowing easy scalability and flexibility in handling numerous sensors and devices.
-  • QoS levels and reliability: MQTT’s different QoS levels enhance reliability by ensuring message delivery according to the application’s criticality.
-QoS 0 delivers messages at most once without acknowledgment.
-QoS 1 ensures messages are delivered at least once, while QoS 2 guarantees exactly one delivery through a four-step handshake, preventing duplication and ensuring data integrity for critical communications.
-Moreover, MQTT’s LWT (Last Will and Testament) feature allows the system to detect and respond to unexpected device disconnections, maintaining the robustness of the network.
-4 Proposed DT pipeline
-Figure 2: Architecture of the proposed DT pipeline: Urban Hybrid Twin.
-In this section, we present a DT architecture called “Urban Hybrid Twin,” based on the sensing/communication/computing testbed deployed in NYC (Fig.
-1).
-The proposed architecture of “Urban Hybrid Twin” is enabled by CPS, user software defined radios (SDR), cameras and LiDARs, high speed communications, and AI computing deployed in the edge cloud.
-Figure 3: Diagram of intersection safety warning use case.
-4.1 System components
-4.1.1 Physical Infrastructure
-The physical functionality of the proposed DT is built on top of the COSMOS testbed (“Cloud enhanced Open Software defined mobile wireless testbed for city-Scale deployment” [125] developed for real-world research, development, and deployment of city-scale advanced wireless communications and innovative applications [125].
-It targets the technology “sweet spot” of ultra-high bandwidth and ultra-low latency, a capability that will enable a broad class of safety-time critical applications.
-Deployed in West Harlem, NYC, next to Columbia University campus, the COSMOS testbed is an enabler for research on the design, development, and deployment of a DT for urban traffic management.
-Pilot DT experiments are executed at the intersection of Amsterdam Avenue and 120th Street, where COSMOS cameras, LiDAR, and wireless sensing and communication nodes are deployed.
-The cameras employ H264 encoding with an I-frame interval set to 10 frames.
-The live RTSP video streams–with 4K resolution at 30 fps–were processed on a COSMOS edge server equipped with an A100 GPU.
-The LiDAR is a Velodyne VLP-32C with up to 200m range.
-Note that there is a growing trend of studies relying on emerging communication protocols like C-V2X for vehicle DTs [145, 157], which requires customized infrastructure and user-end devices.
-Our initial experiments utilize contemporary communications and sensing protocols for rapid deployment.
-They will be replaced by emerging methods that can achieve lower latency and higher bandwidths.
-Software defined features of the COSMOS testbed and the flexibility of radios make it possible to deploy emerging technologies such as C-V2X and experiment with novel low-latency applications.
-4.1.2 Digital twin
-Building on data collected from the physical world, we built an Unreal Engine based simulation platform for generating photorealistic street scenes.
-Moving objects are populated into an integrated SUMO-CARLA simulation platform to validate our safety warning application [43], where vehicle locations are synchronized from the real-world MQTT messages.
-4.2 System functionality
-4.2.1 Resource optimization
+Intersections, where sixty percent of crashes happen [135, 136], are critical bottlenecks of an urban transportation network.
+To improve urban road safety and increase traffic capacities, safety warning is the key.
+Leveraging existing sensors at the intersection, we have developed a combined VRU and vehicle warning application.
+The cameras and the LiDAR sense the presence of VRUs at urban intersections, make predictions of their movements, apply traffic operation and control strategies, and feedback to system controllers and road users, with the primary goal of increasing traffic safety and efficiency (see Fig.
+3).
+Resource optimization
 Due to the computational needs of complex computer vision models and the high volume of video data, ensuring scalability in a video analytics pipeline requires optimization of its configuration to maintain performance.
@@
 2) component that continuously adapts the system’s configuration parameters to maintain its performance.
-4.2.2 Use cases
-4.2.2.1   Pedestrian-vehicle safety warning
-Leveraging existing sensors at the intersection, we have developed a combined VRU and vehicle warning application.
-The cameras and the LiDAR sense the presence of VRUs at urban intersections, make predictions of their movements, apply traffic operation and control strategies, and feedback to system controllers and road users, with the primary goal of increasing traffic safety and efficiency.
-(see Fig.
-3).
-A variety of AI models have been developed to predict future trajectories of interacting road users at the intersection, including InfoSTGCAN (An Information-Maximizing Spatial-Temporal Graph Convolutional Attention Network) integrating InfoGAN and spatiotemporal graph convolutional attention network [131], and PI-NeuGODE using physics-informed graph neural ordinary differential equations[107, 110], and uncertainty quantification (UQ) to characterize the predictive confidence [105].
-4.2.2.2   Learn to optimize adaptive traffic signal control
-To optimize traffic signal control over a road network requires first to predict traffic patterns on road segments and intersections.
-The AI models we developed include PINO (physics-informed neural operator) [45] that predicts future traffic dynamics by training a mapping from initial conditions to traffic states in time and space, building on our previous work on PIDL [138, 139] and the UQ problem [108, 106].
-For ATSC, we have developed an RL method, CVLight [109], and an FRL method [42] on the COSMOS network [42].
-Their performances are illustrated in Fig.
-4.
-Figure 4: (a): ATSC for a traffic corridor. (b): Performance metrics of ATSC methods. Left: travel time and communication cost; Right: traffic volume on the incoming lane over time.
-4.3 Evaluation
+In addition to dynamic adaptation of the system’s configuration, we can reduce latency and GPU consumption by identifying Regions of Interest (RoIs) where pedestrians might be in danger, using lightweight processing (e.g., low resolution, small models).
+As shown in Table VIII, smaller models (“YOLOv8s” in Column 4) incur significantly lower latency.
+Larger models (“YOLOv8x” in Column 6) with higher resolution and frame rate are only triggered when a critical danger area, i.e., RoI, is detected.
+For example, detecting large objects such as vehicles does not require large models or high resolution.
+Therefore, we can use a smaller model and lower resolution to detect vehicles and their trajectories, and based on that, determine the danger areas where pedestrians may be at risk or in the blind zones of vehicles.
+These identified danger areas are then processed using larger models and higher resolution to detect pedestrians at risk and notify them or the vehicles if necessary.
+In our intersection safety warning system (shown in Fig.
+3), we have embedded an RoI Specifier element to facilitate this approach.
+AI models
+A variety of AI models are developed to predict future trajectories of interacting road users at the intersection, including InfoSTGCAN (An Information-Maximizing Spatial-Temporal Graph Convolutional Attention Network) [137], which encodes road user trajectories into quantized latent codes to account for heterogeneity in road users, PI-NeuGODE (Physics-Informed Graph Neural Ordinary Differential Equations), using physics-informed graph neural ordinary differential equations [138, 139], and uncertainty quantification (UQ) to characterize the predictive confidence [140].
+These probabilistic trajectory prediction methods enhance pedestrian safety at intersections by incorporating UQ into risk assessment.
+The Kalman Filter (KF) [141] performs prediction by recursively estimating the state of a moving object using a linear dynamic model with Gaussian noise.
+Trajectron++ [142] models multimodal future trajectories by combining recurrent neural networks with conditional variational autoencoders, enabling probabilistic and socially-aware predictions.
+We present the predicted trajectories and performance comparison in Fig.
+7, where the performance is evaluated using Average Displacement Error (ADE, i.e., the mean Euclidean distance between predicted and ground-truth positions over all future time steps) and Final Displacement Error (FDE, i.e., the Euclidean distance at the final prediction step).
+Method ADE $\downarrow$ FDE $\downarrow$
+Kalman Filter 0.91 1.92
+Trajectron++ 0.77 1.34
+ADE and FDE comparison. Lower values are better ($\downarrow$).
+Figure 7: Performance comparison for trajectory prediction
+Evaluation
 In safety-critical applications, it is infeasible to perform field experiments for such assessment. DT simulation is thus used to assess the effectiveness of the system.
-4.3.1 End-to-end latency
-Measuring latency could be particularly problematic due to the need for cross-network timing synchronization of devices and compute server.
-Spot measurements of several components are shown in Table 6.
-This table presents the latency incurred by the main components of a typical video analytics pipeline for object detection and tracking, as depicted in Fig.
-3, which has been systematically optimized in terms of memory and resource usage.
+a. End-to-end latency
+Measuring latency could be particularly challenging due to the need for cross-network timing synchronization of devices and compute server.
+Table VIII presents the latency incurred by the main components of a typical video analytics pipeline for object detection and tracking, as shown in Fig.
+3.
+We measured the latency for three sizes of YOLOv8 object detection model: YOLOv8s (small, $\sim$11.2 million parameters), YOLOv8m (medium, $\sim$25.9 million parameters), and YOLOv8x (large, $\sim$68.2 million parameters).
+Larger models offer improved detection accuracy but come at the cost of increased latency and higher GPU and memory usage [143].
+The pipeline has been systematically optimized in terms of memory and resource usage.
 These elements usually run on an edge device or server.
-The results in Table 6 identify potential bottlenecks within the pipeline and indicate which components could benefit from further optimization.
-Additionally, the table illustrates the impact of video content complexity on the latency of each component, providing insights into performance variations during busy and sparse traffic conditions.
-To understand the sensitivity and complexities of real-time video analytics, we measured the latency introduced by key components of a typical video analytics pipeline, illustrated in Fig.
-3.
-For object detection, we used the YOLOv8m model [149] (batch size: 1) with the Nvidia DCF-based tracker.
-The YOLOv8m model was configured to accept input images at 832×832 resolution, corresponding to the size of the regions cropped by the RoI cropping component.
-The 5G and LTE networking latency was measured using iPhone 13 Pro Max, while wired and wireless networks provided by Columbia University Information Technology were used for the Ethernet and WiFi measurements.
-Latency measurements are conducted at various times of the day with different traffic densities.
-Tab.
-6 summarizes the results of these measurements by comparing the average and standard deviation of latency per frame for the main components of the pipeline.
+The results identify potential bottlenecks within the pipeline and indicate which components could benefit from further optimization.
+TABLE VIII: Average latency and Standard Deviation for different pipeline elements.
 Metric Inference Elements Downstream Elements
-Reception Pre-processing Object Detection Object Tracking MQTT Msg Creator MQTT Msg Retrieval
-Ethernet Wi-Fi LTE 5G
-Avg Latency (Std Dev) [ms] 1.94 (1.69) 0.108 (0.024) 7.216 (0.121) 0.973 (0.173) 0.081 (0.021) 3.21 (0.315) 6.86 (1.19) 45.72 (15.30) 39.21 (7.12)
-Table 6: Average Latency and Standard Deviation for different pipeline elements.
-These latency measurements demonstrate how each pipeline component influences overall latency and its variability under different video content densities.
-The results highlight the necessity of optimizing these components for real-time applications, particularly in environments with dynamic traffic patterns.
-In the following, we describe the function of each element in the inference pipeline (see Fig.
-3) and their associated latencies based on the results presented in Tab.
-6.
-  • Reception: Receives RTP (Real-time Transport Protocol) data packets and, after parsing and decoding, converts them to raw video frames.
-The latency increases under busy traffic conditions, primarily due to higher complexity of the data being processed, such as larger frame sizes and more variation in the number of data packets per frame.
-  • Pre-processing: Extracts regions of interest (RoIs) from raw video frames and discards irrelevant regions.
-It incurs low latency with minimal variations, indicating the effectiveness of pre-processing in optimizing video analytics performance.
-  • Object detection: Applies the object detection model to the cropped RoIs.
-This element incurs the highest latency in the pipeline.
-The latency is slightly higher in busy traffic/crowd conditions due to more objects in each video frame.
-  • Object tracking: Assigns unique IDs to detected objects across video frames for tracking.
-It has higher latency under busy conditions due to the increased number of objects being tracked simultaneously.
-  • MQTT message creator: Converts the extracted metadata into MQTT format [104] and transmits it over a designated topic.
-Its latency is consistently low, with a slight increase during busy conditions due to the need to assemble larger messages.
-  • MQTT message retriever: A web application captures round-trip time (RTT) of MQTT messages, and measures the latency as half an RTT.
-4.3.2 Safety warning performance
+Reception Pre-processing Object Detection Model Object Tracking MQTT Msg Creator MQTT Msg Retrieval
+Small (YOLOv8s) Medium (YOLOv8m) Large (YOLOv8x) Ethernet Wi-Fi LTE 5G
+Avg Latency (Std Dev) [ms] 1.94 (1.69) 0.108 (0.024) 4.034 (0.084) 7.216 (0.086) 11.140 (1.800) 0.973 (0.173) 0.081 (0.021) 3.21 (0.315) 6.86 (1.19) 45.72 (15.30) 39.21 (7.12)
+b. Safety message generation accuracy
 Figure 8: A simulated pedestrian in CARLA (left) surrounded by cars with real-world positions (right).
-To validate the accuracy of our trajectory prediction and risk assessment algorithms for warning generation, we conducted three rounds of simulation, each lasting $1010$ minutes and generating a total of $232232$ virtual pedestrians in CARLA.
+To validate the accuracy of our trajectory prediction and risk assessment algorithms for warning generation, we conducted three rounds of simulation, each lasting $10$ minutes and generating a total of $232$ virtual pedestrians in CARLA.
 We then compared the number of collision warning messages with the actual number of simulated collisions.
 The process is illustrated in Fig.
-5.
-Metrics for risk assessment include time to collision (TTC) (i.e., the time remaining before a collision occurs) and post-encroachment time (PET) (i.e., the time interval between when the encroaching vehicle leaves the conflict point and when the vehicle of the right-of-way arrives at the conflict point).
-For prediction, we select TTC and set its threshold as $1.21.2$ seconds to minimize both false positives (FP) and false negatives (FN) while optimizing true positives (TP) and true negatives (TN).
-The resultant confusion matrix is $\text{TP},\text{FP}],[\text{FN},\text{TN}]]=[[66,45],[2,119]].[ [ TP , FP ] , [ FN , TN ] ] = [ [ 66 , 45 ] , [ 2 , 119$ ] ] .
-5 Conclusions and open questions
-Transportation applications in urban settings are generally challenging to design, develop, deploy, and test for their potential unsafe and unethical consequences.
-Thus, AI-empowered DT plays a critical role in effective implementation of these applications.
-Despite relatively sparse literature in this domain, we review an ensemble body of literature on how to leverage emerging technologies in sensing, communication, edge and cloud computing, for urban traffic management.
-We hope this paper can serve as a pointer to help researchers and practitioners identify challenges and opportunities on the development of DTs; a bridge to initiate conversations across interdisciplinary researchers in sensing, communication and networking, traffic engineering, and AI; a road map to exploiting potentials of DTs for diverse transportation applications.
-5.1 Emerging trends
+8.
+(a) TTC (in seconds)
+(b) Distance (in pixels)
+Figure 9: ROC curves for trajectory prediction to determine optimal thresholds
+Note that metrics for risk assessment include time to collision (TTC) (i.e., the time remaining before a collision occurs) and post-encroachment time (PET) (i.e., the time interval between when the encroaching vehicle leaves the conflict point and when the vehicle of the right-of-way arrives at the conflict point).
+TTC is commonly used, while PET is typically for post-event analysis.
+To select the optimal thresholds of TTC, we first generate the ROC (Receiver Operating Characteristic) curve in Fig.
+9.
+Fig.
+9(a) illustrates the relationship between the true positive rate (TPR) and the false positive rate (FPR) across various TTC threshold values, ranging from 0.1 to 1.2 seconds.
+The optimal TTC threshold is 1.1 seconds (indicated by a blue square), at which the TPR reaches $0.958$ and the FPR is $0.4$.
+To compute TTC, we need to compare each pair of predicted trajectory points in the next $t$ time steps, determined by a danger distance threshold that is defined as when the proximity between a vehicle and a pedestrian constitutes a dangerous interaction.
+Fig.
+9(b) presents the TPR and FPR for the threshold of danger distances ranging from $5$ to $100$ pixels.
+The optimal threshold is valued at $30$ pixels, which yields a TPR of $0.958$ and an FPR of $0.345$.
+Based on the above two thresholds, we run the collision prediction model.
+The resultant confusion matrix for predicted collisions under the selected thresholds is given by $\text{TP},\text{FP}],[\text{FN},\text{TN}]]=[[66,45],[2,119]]$.
+c. Human response time assessment
+Figure 10: Response time determination in the VR experiment
+Figure 11: Survey results from the VR experiment
+Would issuing warnings to pedestrians help reduce users’ response time and increase their safety awareness?
+To test this hypothesis, we designed virtual reality (VR) experiments in Unity3D [144], where warnings are provided via voice and text displayed on the VR headset Meta Quest 3.
+There are two traffic scenarios, one involving an interaction between the participant (i.e., a pedestrian) and an oncoming scooter, and the other between the participant and an oncoming vehicle.
+In each scenario, participants first receive no warning and then a warning, and their response times are recorded.
+As illustrated in Fig.
+10, when traffic approaches at time $t$, the user either receives a warning or does not.
+The user stops after a response time $t_{r}$, which is then recorded.
+The response time distributions are presented in Fig.
+12, with the ‘no warning’ condition shown in yellow and the ‘warning’ condition in blue.
+The issuance of warning has reduced people’s average response time by $0.62\mathrm{s}$ (in Scenario 1) and $1.11\mathrm{s}$ (in Scenario 2), respectively.
+The demographics of the participants are shown in the upper section of Fig.
+11.
+The bar chart summarizes the participants’ survey responses regarding VR and traffic realism, the perceived importance of safety warnings, and comfort within the VR environment, with mean and standard deviation indicated on the top left corners of each subfigure.
+Figure 12: Comparison of response times with and without a warning. A scooter (in Scenario 1) and a vehicle (in Scenario 2) poses the danger, respectively.
+IV-C3 Discussions
+We would like to highlight that our NYC’s testbed deployment demonstrates generalizability.
+As the biggest metropolis of the U.S., NYC’s dense population, limited space, and multimodal transportation systems (consisting of public transit with buses and subway, driving, ferry) motivate us to study the interactions between automobiles and VRUs (including pedestrians, cyclists, and surging micromoblity and e-mobility).
+Located in one of the world’s busiest regions - Manhattan of the NYC, the COSMOS testbed offers a natural laboratory for us to collect abundant data and conduct live experiments, which would otherwise not be possible.
+We leverage the COSMOS testbed to design and implement sensor (e.g., camera, LiDAR) data analytics systems and evaluate them in urban settings to ensure they can scale to thousands of connected intersections.
+To achieve this, we ensure that the designed analytics systems are efficient in resource consumption, particularly network bandwidth and GPU usage.
+We also ensure that these systems are modular and deployable in a distributed manner across multiple edge servers and cloud infrastructure.
+Fig.
+3 demonstrates such a modular and distributable system.
+This design enables flexibility, allowing the system to be deployed in settings where a single edge server cannot handle the full workload, requiring distribution across several edge servers.
+Therefore, the system must be designed such that each module can be easily deployed on a different server and can seamlessly communicate with other modules, regardless of whether they reside on the same server.
+Additionally, the AI models used must be replaceable or generalizable to accommodate different environments, camera angles, heights, and other variations.
+With these principles in mind, we use the COSMOS testbed to develop scalable systems.
+The NYC’s dense high-rise buildings enable us to leverage legacy infrastructure for traffic sensing and monitoring, and facilitates the instrumentation of cameras and communication nodes.
+Note that there is a growing trend of studies relying on emerging communication protocols like C-V2X for vehicle DTs [145, 146], which requires customized infrastructure and user-end devices.
+For rapid and scalable deployment, our pilot experiments utilize contemporary protocols, which hold the potential to further leverage the city’s legacy traffic cameras and wireless communication infrastructure for scalability and cost-effectiveness.
+Our technologies could be replaced by emerging methods that can achieve lower latency and higher bandwidths.
+For instance, software defined features of the COSMOS testbed and the flexibility of radios make it possible to deploy emerging technologies such as C-V2X and experiment with novel low-latency applications.
+Since major intersections in the neighborhood has similar sensor suites, it would not need additional infrastructure deployment to generalize our system to network wide multi-intersection settings.
+We will extend our use cases to multi-pedestrian warning scenarios.
+Smartphone‑to‑cloud beaconing with an Adaptive Multi‑Mode (AMM) scheme enables real‑time location sharing from multiple devices, allowing the server to issue timely risk alerts [147].
+A Vehicle‑to‑Pedestrian (V2P) system [148] leveraging Bluetooth Low Energy (BLE) advertising on smartphones broadcasts standardized Personal Safety Messages (PSM) to address the multi‑device communication challenge in road safety.
+Moreover, efficient multi-device communication is essential for real-time data sharing and coordination among mobile users in smart urban environments.
+Key challenges include ensuring reliable connectivity in dense settings, minimizing communication latency, and maintaining user privacy.
+Advanced techniques such as device-to-device (D2D) communication [149], opportunistic networking [150], and privacy-preserving data aggregation [151] have been developed to address these issues.
+V Conclusions and open questions
+In this paper, we first review the AI methods applied to every stage of the DT pipeline, from object detection, tracking, prediction, simulation, to traffic operation and management.
+Then, leveraging the unique characteristics of NYC, we propose a DT architecture and present a safety-critical use case, namely, intersection safety warning.
+Three evaluation methods are performed, including measuring latency, simulating collision reduction, and VR experiments for human responses.
+A growing number of sensors, explosive amounts of data, and increasing computational powers have opened up tremendous opportunities for researchers to apply AI to create, train, evaluate, and streamline DT pipelines for uban traffic management.
+Subsequently, we will present emerging trends, challenges, and open questions for the development of DT.
+V-A Emerging trends
 While literature on DTs for individual components has been surging, how each element of the DT works collectively and function organically is key to the development of the next-generation DT, which could empower the intelligence and automation of transportation applications.
-5.1.1 Engineering the pipeline
+V-A1 Engineering the pipeline
 Engineering a DT via the integration of multiple subsystems poses technical challenges, and we will name a few.
-5.1.1.1   Datasets for urban computer vision
-Due to limited availability and single perspective of real-world datasets, synthetic data generation from realistic 3D simulators has become increasingly popular for urban computer vision problems.
-The SYNTHIA dataset, for example, provides a diverse set of images from a simulated city with pixel-level semantic annotations, aiding tasks like semantic segmentation and scene understanding [129].
-CARLA has been widely used in object detection and segmentation [115, 59, 101].
-Similarly, GTA V has been employed to collect game frames for training autonomous agents [128].
-Unreal Engine has gained traction as a rendering engine for various vision tasks [4, 55, 29].
-MatrixCity utilizes Unreal Engine 5’s City Sample project to benchmark neural rendering models, focusing on pedestrian- and vehicle-free environments [78].
-The Boundless project shows YOLOv8x’s enhanced performance on real-world datasets when trained with Unreal Engine 5-synthesized data [151].
-However, mixed-perspective datasets that adapt to diverse deployment conditions remain scarce.
-Advances in open-vocabulary object detection offer a promising avenue for addressing this gap [92, 103, 24].
-5.1.1.2   Sensor fusion with time synchronization
+I. Sensor fusion with time synchronization
 Various methods have been proposed to tackle the time-synchronization problem in multi-camera settings for a single intersection or area, either utilizing visual cues or through explicit clock synchronization between the cameras.
@@
 Efficient, low-latency implementation of these methods are vital for synchronization and fusion of camera predictions.
-5.1.2 Designing edge-cloud architecture and networking
+II. Designing edge-cloud architecture and networking
 To facilitate the development of a DT in an urban setting with numerous intersections capable of communicating with vehicles and VRUs, an extended network of cloud-connected edge devices and sensors, such as cameras, is required.
@@
 To this end, data and device federation using federated (reinforcement) learning has gained growing traction [162, 163, 164].
+III. Integrated sensing and communication
 Integrated sensing and communication (ISAC) is an emerging direction in the design of Beyond-5G wireless networks, enabling transmitted communication waveforms to be opportunistically used as radar-like sensors [165, 166].
@@
 Future research directions may focus on leveraging high angular resolution from densely packed phased arrays to achieve precise beamforming, with the potential to introduce imaging-like capabilities within communication networks [167, 168].
-5.1.3 Physics-informed DT development
-The mutual interaction between domain knowledge and AI is necessary.
+V-A2 Emerging DT applications
+With the increasing demand for urban passenger and goods delivery, emerging technology like urban sensing, electrification, connectivity and autonomy, and robotics have been gradually transforming urban streetscapes, which pose new challenges to the operation and management of urban infrastructure and public space.
+DTs are crucial to improve urban safety and mobility in infrastructure planning, service operation and management, and ultimately, policy making.
+Here we do not aim to enumerate comprehensive urban transportation applications with DTs, since there exist survey papers [13, 14, 4, 5, 6], which summarize those in operation (e.g., anomaly detection and warning, emergency response), maintenance, and mobility (e.g., transit operation, or driving).
+Instead, we focus on emerging applications in urban settings, and the potentials the DT holds for them.
+In Fig.
+13, emerging use cases are categorized based on required communication latency (x-axis), spatial resolution (y-axis), and data bandwidth (z-axis), respectively.
+Figure 13: Urban T-DT use cases.
+Stochasticity arising from travel demands, un-predicted VRU movement, and traffic gridlocks requires traffic operators to design adaptive traffic signal control (ATSC).
+With a large amount of agents (including VRUs, traffic lights) continuously interacting in a stochastic environment, classical optimization tools in deterministic environments could fail to capture such complex decision-making processes.
+There is a surge in employing learning methods, including reinforcement learning and federated learning, to optimize traffic signals [169, 170, 171, 172, 173, 174, 175, 176, 177].
+With an increasing concern in data governance and privacy, federated learning witnesses a growing trend that allows optimizing centralized control while preserving distributed data privacy [178, 179, 180, 164, 181, 182].
+Driven by growing urban populations, low-altitude economy aims to exploit vertical space in urban environments, advanced by drones and eVTOL (electric vertical take-off and landing).
+It will foster wide applications, ranging from urban air mobility, urban logistics, agriculture, emergency service, to infrastructure inspection, surveillance and security.
+Critical challenges arise in sensing, modeling, predicting urban aerodynamics, and collision avoidance with high-rise buildings and flying objects.
+Open questions include market demand estimation, infrastructure planning, operation and design.
+DTs hold the potential to the adoption of these emerging technologies, for its power in simulating emerging demand patterns, modeling their kinetics and dynamics, routing and charging behaviors.
+To build such DTs, CPS technologies and AI methodologies would facilitate precise sensing and perception [146], reliable and fast communication and networking [183], seamless multimodal coordination (such as drone-truck-mobile robot delivery) [184], as well as safe and optimal decisions [185].
+V-A3 AI models
+AI methods have been applied to every stage of the DT pipeline. We will point out key challenges of AI methods in transportation.
+Physics-informed AI
+When it comes to the development of a T-DT, the mutual interaction between domain knowledge and AI is necessary.
 Domain knowledge in traffic flow and road safety that has been developed for decades, provides valuable insights into every stage of the DT pipeline.
@@
 For instance, the understanding of how traffic evolves across time and space and how entities interact with one another at a road junction, could guide to what degree of granularity and precision the semantic segmentation should be done in object detection and tracking, how communication and networking resources should be prioritized, and what models and metrics should be used for risk prediction.
-5.1.4 Safety guarantee of AI
+Safety guarantee of AI
 DTs, heavily relying on hardware, software, and algorithms, are vulnerable to risks posed by probabilistic events.
 Failures in sensors, communication channels, or algorithms can result in erroneous or missing inputs to downstream prediction or decision-making, ultimately compromising the feedback loop from the DT to the physical.
-The rise of AI-empowered DTs introduces additional risks, including adversarial attacks, and trustworthiness in real-world applications.
+The rise of AI-empowered DTs introduces additional risks, including adversarial attacks [186, 187, 188], ethical concerns [189, 190, 191], and trustworthiness in real-world applications [192, 193, 194].
 How can DTs be designed with provable safety guarantees, accounting for uncertainties, failures, and attacks?
@@
 Additionally, control barrier functions and Lyapunov functions have been integrated into neural networks to enforce safety constraints and maintain system stability [203, 204].
-5.1.5 Sim-to-real gap closing
-What distinguishes a traffic DT from a conventional traffic simulator?
-Capabilities of generating simulations and predictions consistent with the real world is the key.
-There is a growing amount of studies focused on sim-to-real policy learning [196, 15, 62], i.e., how optimal policies learned from a DT can be generalized to real-world implementation.
-The larger the sim-to-real gap, the less reliable the DT is.
-To the best of our knowledge, there is little research on how to close the sim-to-real gap leveraging transportation DTs, which we believe is the key to transportation applications.
-5.2 Open questions: Prototyping DTs for human intelligence
+Evaluation and validation
+As opposed to traditional traffic management systems heavily involved with humans, an AI-powered CPS-enabled traffic management system demands high degree of autonomy, with increased complexity and scale.
+Since testing traffic management strategies could be unethical and unsafe, DT thus becomes a crucial tool for the test and verification.
+There does not exist a unified scheme about what to validate, verify, and test in a DT.
+Here, we would like to decompose this problem into several layers.
+First, we need to evaluate whether an DT represents its corresponding physical world correctly.
+Ideally, the digital is expected to be the twin of the physical, accordingly, “Grieves performance test” [1] is a high-level abstract way to compare the difference in the outputs of both the physical and the DT.
+This is associated with real2sim gap to be defined and elaborated more in the next section.
+Second, we need to verify the system behavior of an DT-enabled CPS, and ensure that it performs as desired [205], such as safety, efficiency, accuracy, and timeliness.
+Such testing is non-trivial, due to the integration of cyber and physical components, as well as their two-way coupling via communication and networking.
+There are normally four types of tests [205], namely, conformance testing (i.e., whether a system conforms to an expected behavior), robustness testing (i.e., whether a system is robust against stochasticity in environments), fragility testing (i.e., whether the output of a system is robust against perturbation in inputs), and security testing (i.e., whether a system is not affected by cyber-attacks).
+DTs play important roles in the above tests, while experimental design is crucial to cover comprehensive scenarios leveraging game theory [206] and in recent years, generative models [207].
+The presence of human factors along the DT pipeline could complicates the evaluation of cyber-physical-human systems, because of randomness and unpredictability in human behaviors [208].
+Humans are not only participants in traffic (as drivers, pedestrians, cyclists), they are also creators and users of DTs.
+For DTs that involve the feedback to humans, human-in-the-loop test is widely used [209, 210, 208].
+Recent years have witnessed a growing trend of using augmented reality and virtual reality to engage humans as pedestrians, cyclists, or scooters in virtual environments without inducing real-world risks [134, 211].
+Hardware-in-the-loop testing, such as vehicle-in-the-loop [212], could help test the performance of some physical components “live,” while allowing the rest to be simulated within a DT.
+Since there does not exist a unified approach for the DT testing, evaluation methods, metrics, testing platforms should be developed to facilitate standardized assessment and validation of AI models.
+Benchmark datasets and methods
+To advance the application of AI in DT, we must “stand on the shoulders of giants.” In other words, each team does not simply develop one DT for its own internal use.
+Instead, we hope that an AI-powered DT could be generalized to diverse tasks, transferred to diverse spatiotemporal settings, and shared for co-development among global researchers.
+Accordingly, we need to standardize application scenarios, benchmark datasets and methods, and unified test environments, for repetitive training and test [213].
+Benchmark datasets and methods necessitate performance comparison of any newly proposed AI methods against the state-of-the-art (SOTA) methods.
+Thus, the transportation community must push to open source data, codes, simulation, algorithms, and results for replicability.
+For example, standard test platforms, such as Gym [214], Flower [215], PettingZoo [216, 217], and Ray RLlib [218], are important to benchmark various AI models and methods, which is mostly missing in the transportation community.
+Moreover, open mixed-perspective datasets that adapt to diverse deployment conditions remain scarce.
+Advances in open-vocabulary object detection offer a promising avenue to address this gap [219, 220, 221].
+V-B Challenges and open questions
+V-B1 Closing real-to-sim-to-real gap
+A key challenge in applying models trained in simulated environments to the real-world is domain shift between the training and test environments.
+Models trained in simulators might likely experience performance degradation when tested in the real world, where the environment includes unpredictable variations that simulators cannot fully replicate.
+Such a shift – caused by discrepancies between simulated and real-world conditions – can undermine the generalization of models, resulting in reduced performance in scenarios not represented in training.
+What distinguishes an DT from a conventional computer simulator?
+We believe the key lies in its capability of generating simulations and predictions consistent with the real world, as well as preserving intervention performance.
+Accordingly, two gaps exist while establishing a DT, namely, real-to-sim (real2sim) gap (i.e., the deviation of the simulated digital world from the real-world), and sim-to-real (sim2real) gap (i.e., the performance deviation of interventions implemented in the real-world from those simulated in the digital world).
+The smaller these two gaps are, the closer an DT is to reality.
+Although these gaps penetrate through each stage along the DT pipeline, from computer vision, tracking, to prediction, and intervention, sim2real transfer is more studied in object detection [222] and policy learning [223, 224, 225, 226].
+The major application area of sim2real transfer in DT is robotic manipulation [227], while that in urban navigation and autonomous driving has witnessed a gradual surge, participially in computer vision [228] and reinforcement learning [229, 230].
+Key methods include domain randomization [231, 207], which introduces variations to augment training data and expose the model to a wide range of scenarios; and domain adaptation [232, 233], a technique of finding mappings to transfer data points observed empirically from two different data distributions.
+How do we characterize real2sim gap and control such a gap?
+This boils down to quantifying errors of the digital representation of a physical world.
+Minimizing the real2sim gap is key to system identification and representation learning.
+Depending on observability and internal workings, a system could be modeled as white-box, grey-box, or black-box [234].
+Domain knowledge could help represent the physical world with higher accuracy.
+For example, hybrid twin [235, 236] relies on both data-driven and physics-informed AI [237].
+Since calibration of a full model could be time-consuming, expensive, and potentially infeasible, the model reduction philosophy has become popular.
+Digital shadow [11], a digital model with one-way data exchange from the physical to the digital, takes less effort to build, but could fail to update the state of the physical world once feedback is executed.
+Digital cousin [238] aims not to build a simulation model that replicates the reality exactly, instead, mainly focus on end-to-end gap, namely, from real-world sensing to intervention.
+V-B2 Prototyping DTs for human intelligence
 Many studies define a DT as emerging technologies, namely, object detection and tracking with edge-cloud computing and communication.
-These technological enablers, however, are actually the “eyes” of a DT, while what really distinguishes a DT from traditional simulators lies in its “brain,” the prediction and decision making modules that are capable of extracting patterns, modeling semantics, and making informed decisions drawing upon what has been seen and perceived.
-Ultimately, we should produce a next-generation of DT ideally with human-like intelligence but automation capabilities.
-A DT with comparable human-like intelligence or artificial artificial general intelligence (AGI) should consist of a hierarchical cognitive structure analogous to a human’s biological system, backed up by emerging hardware, software, AI algorithms, and API interfaces.
-The architecture of a proposed DT could consist of (see Fig.
-6):
+These technological enablers, however, are essentially “eyes” of a DT, while what really distinguishes a DT from traditional simulators lies in its “brain,” the prediction and decision making modules that are capable of extracting patterns, modeling semantics, and making informed decisions drawing upon what has been seen and perceived.
+How do we establish a foundational DT that develops human-like intelligence with machine automation?
+A DT with comparable human-like intelligence or artificial general intelligence (AGI) should consist of a hierarchical cognitive structure analogous to a human’s neural system, backed up by emerging hardware, software, AI algorithms, and API interfaces.
+The architecture of a proposed DT could consist of:
   1. 1.
@@
   2. 2.
-    Brain: data storage and processing, reasoning and planning, inference and generalization, powered by AI;
+    Neural systems: edge-cloud networking and computing backbone, powered by resource allocators [239] and cognitive DT [240];
   3. 3.
-    Neural systems: edge-cloud networking and computing backbone, powered by resource allocators;
+    Brain: data storage and processing, reasoning and planning, inference and generalization, powered by causal inference and counterfactual analysis [241];
   4. 4.
-    Communication: natural language processing, powered by large language models (LLM) and foundation models.
-Figure 6: Schematic diagram of a DT analogous to humans.
-A growing number of sensors and explosive amounts of data have opened up tremendous opportunities for researchers to apply AI to create, train, evaluate, and streamline traffic DT pipelines, from object detection, tracking, prediction, simulation, to traffic operation and management.
-The emerging LLM and foundational models have further demonstrated great potential in DTs.
-LLMs could generate new data for training data augmentation or privacy concerns [165, 191, 21], enhance interactions between human users and the DT system [188], making personalized recommendations [177, 38] and even automate the creation process of DTs [10].
+    Communication and reasoning: natural language processing and vision reasoning, powered by generative AI (GenAI) [242].
+A T-DT could be deemed as the world model of a transportation manager.
+A world model is a mental model learned by an AI agent to simulate the evolution of its environment for action planning and reasoning.
+It is a special type of DT that relies on the agent’s own sensor information.
+Moreover, it can be embedded into a DT as the AI agent’s internal, abstract representation of the physical world.
+A world model emerges from the field of robotic learning, and there is an emerging trend to augment a world model with cognitive and reasoning capabilities for more accurate representation and prediction [243].
+Such a trend, we strongly believe, must be the pathway for the next-generation of DTs, despite that DTs could be an external representation of a system that would facilitate engineers to monitor, diagnose, control, and manage the system.
+In particular, foundational and GenAI models, which are shown to empower cognitive and reasoning architectures of the world model, have demonstrated great potential in DTs [244, 245, 246].
+Large language models (LLM) have been used to generate new data for training [247, 248, 249], enhance interactions between human users and the DT system [250], making personalized recommendations [251, 252] and even automate code generation [253] and creation process of DTs [254].
+On the other hand, vision language models (VLM) help augment training datasets for robustness, including generation of critical events [255], videos [256, 257] and simulations [258], as well as visual question answering [259, 260].
+Thus, we believe that GenAI-powered DT will be the next-generation of DTs for efficient and safe traffic management.
+V-B3 Limits in AI
 Despite the promising future of AI-powered DTs, application of AI to DTs could face challenges.
@@
 Introducing physics based domain knowledge could help fine tune these models, enhance inductive biases, and generate more meaningful outputs.
+On the other hand, the emergence of LLM could facilitate the alignment of AI models with human preference.
+For example, reinforcement learning from human feedback [268] has seen a rapid growth for preferential learning in autonomous driving [269], as well as the generation of more realistic traffic DTs [270].
 Last but not the least, there has been a growing trend in developing safe AI systems aligned with human values and objectives.
@@
 It is thus crucial to continually achieve safety guarantees of AI, especially for the implementation of DTs.
-Acknowledgment
-We would like to thank Mengxuan Liu for her assistance in generating the 3D models.
+In a nutshell, transportation applications in urban settings are generally challenging to design, develop, deploy, and test for their potential unsafe and unethical consequences.
+Thus, AI-empowered DT plays a critical role in effective implementation of these applications.
+Despite relatively sparse literature in this domain, we review an ensemble body of literature on how to leverage emerging technologies in sensing, communication, edge and cloud computing, for urban traffic management.
+We hope this paper can serve as a pointer to help researchers and practitioners understand SOTA methods and gaps on the development of DTs; a bridge to initiate conversations across interdisciplinary researchers; and a road map to exploiting potentials of DTs for urban transportation applications.

### v3

@@
-AI-Powered CPS-Enabled Urban Transportation Digital Twin: Methods and Applications
+AI-Powered CPS-Enabled Vulnerable-User-Aware Urban Transportation Digital Twin: Methods and Applications
 Abstract
@@
 While the majority of studies on the DT focus on its “eyes,” which is the emerging sensing and perception like object detection and tracking, what really distinguishes the DT from a traditional simulator lies in its “brain,” the prediction and decision making capabilities of extracting patterns and making informed decisions from what has been seen and perceived.
-In order to add value to urban transportation management, DTs need to be powered by artificial intelligence and complement with low-latency high-bandwidth sensing and networking technologies, in other words, cyberphysical systems (CPS).
-We will first review the DT pipeline enabled by CPS and propose our DT architecture deployed on a real-world testbed in New York City.
+In order to add value to urban transportation management, DTs need to be powered by artificial intelligence and complement with low-latency high-bandwidth sensing and networking technologies, in other words, cyberphysical systems.
 This paper can be a pointer to help researchers and practitioners identify challenges and opportunities for the development of DTs; a bridge to initiate conversations across disciplines; and a road map to exploiting potentials of DTs for diverse urban transportation applications.
-Index Terms:
-Digital twin, AI, Urban traffic management
-I Introduction
+footnotetext: This work has been published in IEEE Transactions on Intelligent Transportation Systems (IEEE T-ITS).footnotetext: Yongjie Fu, Mehmet K.Turkcan, Zhaobin Mo, and Xuan Di are with the Department of Civil Engineering and Engineering Mechanics, Columbia University.
+Emails: {yf2578, mkt2126, zm2302, sharon.di}@columbia.edufootnotetext: Mehmet K.
+Turkcan and Mahshid Ghasemi contributed equally to this work.footnotetext: Xuan Di, Zoran Kostic, and Gil Zussman are also with the Data Science Institute, Columbia University.footnotetext: Mahshid Ghasemi, Chengbo Zang, Abhishek Adhikari, Zoran Kostic, and Gil Zussman are with the Department of Electrical Engineering, Columbia University.
+Emails: {mg4089, cz2678, abhishek.adhikari, zk2172, gil.zussman}@columbia.edu
+Keywords: Digital twin, AI, Urban traffic management
+1 Introduction
 Urban transportation systems are complex to model and simulate, due to heterogeneous road users (such as cars, pedestrians, cyclists, scooters) interacting in multimodal traffic environments consisting of public and private travel modes.
@@
 With recent explosive growth of literature on DTs, we would like to restrict the scope of this paper to applications in the urban setting, especially when vulnerable road users (VRU) (i.e., non-motorists such as pedestrian, bicyclists, other cyclists, or persons on personal conveyance [199]) are an integral part of the system and also potential users of the DT.
-We will primarily focus on use cases accounting for VRUs with improved traffic safety and efficiency.
 The studies of DTs for urban traffic management, especially involving VRUs, are lacking, partly because the development of a DT for a system is non-trivial, particularly when involved with humans.
@@
 We will depict a DT pipeline prototype, leveraging the architecture of cyberphysical systems and AI methods.
-We will propose a DT for real-time traffic monitoring and optimization, based on an existing physical testbed deployed in New York City (NYC) leveraging cutting-edge sensing, communication, computing, and AI-based automation.
-The overall contributions of this paper include: (1) introducing AI methods used in the DT pipeline; (2) exploring the architecture of transportation DTs and propose a prototype for reference; and (3) identifying gaps and directions.
 The rest of this paper is organized below.
-Section II introduces transportation DTs enabled by cyberphysical systems, and position this paper; Section III reviews the literature along the pipeline of a DT.
-Section IV demonstrates the architecture of our DT, building on a real-world testbed.
-Section V concludes our work, presents potential research directions and open questions.
-II Preliminaries
+Section 2 introduces transportation DTs enabled by cyberphysical systems, and position this paper; Sections 3 and 4 review each component within CPS, and use cases for urban transportation, respectively.
+Section 5 demonstrates the available testbeds for DT testing and validation.
+Section 6 concludes our work, presents potential research directions and open questions.
+2 Preliminaries
 The definition of DT has evolved rapidly.
@@
 Below, we first formalize transportation DT, and then discuss the relation between DT and cyberphyscial systems.
-Definition II.1
+Definition 2.1
 Transportation digital twin (T-DT) is a digital system integrating the pipeline from object detection and tracking, resource allocation, edge-cloud computing and communication, for online simulation, operation, control and management.
@@
 T-DTs are closed-loop with two-way communication, where data, information, and control signals are exchanged with the physical sequentially and reiteratively.
-Definition II.2
-Cyberphyscial systems (CPS) [12] are smart systems that include engineered interacting networks of physical and computational components.
+Definition 2.2
+Cyber-Physical Systems (CPS) [75] are smart systems that include engineered interacting networks of physical and computational components.
 CPS holds great potential to enable real-time applications thanks to emerging technologies in sensing, communication, and computing.
@@
 To enable the technological development of a DT, a physical testbed is needed as a platform for sensing, computing, experimentation, evaluation, as well as design constraints determination.
-In Sec.
-III, we will review the CPS technological enablers needed for the development of a DT, and examine the testbed used for our proposed DT in Sec.
-IV.
-We would like to stress that, this paper aims to discuss how AI algorithms and the architecture of CPS contribute to the urban T-DT (UT-DT) pipeline.
-In contrast, there are highly cited survey papers on T-DT, which are more focused on general transportation scenarios and applications, while AI might not be the key focus.
 Tab.
-I outlines the comparison of a partial set of related work.
-TABLE I: Comparison of survey papers on DT
-Ref. Topic Focus
-[4][5][6] Comprehensive review for vehicle mobility apps Vehicular technology driven pipeline
-[13] Comprehensive review for traffic safety and mobility CPS pipeline highlighting communication
-[14] Comprehensive review for operation and maintenance apps Broad travel modes
-Ours Semi-technical review and position for urban traffic apps AI-Powered CPS pipeline
-We first review various CPS and AI methods needed in the pipeline, and then present our T-DT instance.
-Accordingly, this paper is semi-survey, semi-technical.
-We employ such an organization, because we find that the publications on T-DT are normally segmented by different communities and journals, which could prevent researchers from understanding the entire pipeline, from upstream sensing and perception, to downstream transportation applications.
+1 outlines the comparison of a partial set of related work.
+Table 1: Comparison of relevant review papers on DT
+Ref Topic Focus Urban Applications CPS role in DT VRU Scope Real Testbeds Integration
+Ours Review and position for urban transportation AI-powered CPS architecture Intersection safety warning; Signal control Interlinks physical and cyber layers, DT architecture VRU safety and efficiency A summary of relevant physical testbeds Bidirectional: data exchange, feedback control
+[179] Overview for vehicle mobility Connected and automated vehicles (CAV) CAV design and testing (ADAS, ADS) Cloud-based CPS as methodologies Pedestrian safety in CAV testing Mcity, GoMentum Hardware, software, model-in-the-loop
+[215] Review for vehicle mobility, architecture, case studies AI-based data-driven cloud-edge-device pipeline Personalized ACC, traffic management, driver modeling DT conceptualized as higher-level evolution of CPS Modeled in physical layer of DT Highway 237 in Mountain View, CA Data and service integration across twins
+[89] Review for traffic safety and mobility Hierarchical architecture, cybersecurity Broad transportation safety and mobility services Enables feedback between physical and cyber components As mobile nodes in physical layer, modeled as individual asset for behavior analysis Testbeds mentioned for some literature, but not systematic Integration of data from different levels of twins, and with commun.
+gateway
+[87] Review for smart cities applications Enabling technologies: ML, IoT, CPS, blockchain Transport, water management and automotive technology Facilitates data collection and decision making, DT architecture Ped.
+safety in Ford’s DT for predictive headlights Durham County Council’s streetlights, IIT’s sewage treatment plant, Insurance Institute for Highway Safety test track Interoperability among different DTs within the smart city ecosystem
+[219] Review for transport operation and maintenance Varied levels of adoption across different transport sectors Broad intermodal and ground travel modes CPS enables DT through sensing communication control loop Pedestrian traffic management at train stations N/A Real-time data exchange and synchronization of decision making, integration across platforms
+[223] Review for the full lifecycle of transport infrastructure Functions, architectures, integration along the lifecycle Planning and construction, operation and maintenance, and decommissioning and renewal N/A Integrated into an interoperable system for pedestrian pathways, infrastructure, micro-traffic behaviors N/A Data and system integration across various lifecycle stages
+The publications on T-DT are normally segmented by different communities and journals, which could prevent researchers from understanding the entire pipeline, from upstream sensing and perception, to downstream transportation applications.
 For example, [136] on networking DT is primarily focused on the development of communication technology with evaluation on latency, even its use case is transportation.
@@
 Another example is that, [118] focused on vehicular DT, heavily rely on the foundational knowledge in CPS, which could be somewhat unfamiliar to many transportation researchers.
-Realizing such a gap in a large body of literature on T-DT, this paper aims to unify the knowledge by presenting a comprehensive summary upfront, and exemplify the pipeline following the summary.
+Realizing such a gap in a large body of literature on T-DT, this paper aims to unify the knowledge by presenting a comprehensive summary upfront, and exemplify the pipeline following the summary with selected applications.
 After all, the development of a T-DT calls for the interdisciplinary collaboration across transportation, electrical engineering, mechanical engineering, computer science, and human-machine interaction.
 To foster the readability of the paper, in the next section, we first offer a comprehensive state-of-the-art review on related work along the CPS pipeline.
-III Related work on CPS-enabled DT
+3 CPS-enabled DT
 The development of CPS-enabled DT must engage with expertise in sensing, communication, computing, and human-centric perspectives, where AI methods are backbones.
 Figure 1: Schematic diagram of a DT.
-III-A Sensing and perception
+3.1 Sensing and perception
 Sensors are the “eyes” (and “ears”) of a DT.
 Tab.
-II summarizes the pros and cons of each sensing technology, namely, mobile devices, on-board vehicles, and roadside infrastructure, for urban traffic applications.
+2 summarizes the pros and cons of each sensing technology, namely, mobile devices, on-board vehicles, and roadside infrastructure, for urban traffic applications.
 The former two are mobile sensors with wider spatial coverage but challenge in precision because of moving references, while the latter at fixed locations could face limited sensing ranges and coverage.
-TABLE II: Sensors for urban traffic applications (partially adapted from [16, 17])
+Table 2: Sensors for urban traffic applications (partially adapted from [200, 256])
 Sensor Purpose Advantages Disadvantages
@@
 3. Has difficulty with real-time calibration [247, 248].
-III-B Object detection and tracking
+3.2 Object detection and tracking
 Object detection identifies and classifies objects within the environment using sensors like cameras, Lidar, and Radar.
@@
 Here we thus summarize commonly used and emerging datasets from non-vehicle sensors in Tab.
-III.
-TABLE III: Public datasets for urban traffic scenes (On-board vehicle datasets can be found in [25].)
+3.
+Table 3: Public datasets for urban traffic scenes (On-board vehicle datasets can be found in [45].)
 Sensor location Dataset Purpose Sensor Setup Collection region
@@
 SemanticFormer [190] seeks more structural and humanized environment understanding by constructing a semantic knowledge graph.
-TrajNet++ [82], TDOR [83], and CASPNet++ [84] predict the distributions of future trajectories based on occupancy grid maps.
+TrajNet++ [100], TDOR [80], and CASPNet++ [schäfer2023caspnet] predict the distributions of future trajectories based on occupancy grid maps.
 Models like FRM [160] and PPT [122] decompose the prediction task by taking a multi-stage approach.
@@
 Unlike the tracking algorithms, specific training or fine-tuning is often required before the deployment of trajectory forecasting models to unseen scenarios.
-III-C Real-time video analytics
+3.3 Real-time video analytics
 Developing end-to-end real-time video analytics systems on a large scale for time-sensitive and safety-critical traffic and crowd management applications presents challenges.
@@
 Tab.
-IV provides a comparative analysis of various approaches to address these challenges.
+4 provides a comparative analysis of various approaches to address these challenges.
 The approaches differ in their focus–some prioritize reducing latency and resource consumption, while others emphasize maintaining or enhancing accuracy, especially under constrained conditions.
 This comparison provided in Tab.
-IV highlights the trade-offs inherent in real-time video analytics and emphasizes different optimization methods to balance throughput, accuracy, energy consumption, and computational efficiency across diverse deployment scenarios, including edge devices, cloud platforms, and hybrid environments.
-TABLE IV: Comparison of various approaches and optimization objectives in video analytics.
+4 highlights the trade-offs inherent in real-time video analytics and emphasizes different optimization methods to balance throughput, accuracy, energy consumption, and computational efficiency across diverse deployment scenarios, including edge devices, cloud platforms, and hybrid environments.
+Table 4: Comparison of various approaches and optimization objectives in video analytics.
 References Approach Optimization objective
@@
 Minimize energy consumption for CNN inference services on high-performance GPUs while meeting latency of Service-Level Objectives.
-III-D Communication and networking
-A DT for safety-critical applications requires real time communications with aggressively low latency.
+3.4 Communication and networking
+A DT for safety-critical applications requires real-time communications with aggressively low latency.
 We explore issues related to low latency communications, and survey component technologies and protocols that can be utilized to achieve very low latency.
-III-D1 Real-time requirements and low latency targets
-Sensor and control data in a real time system is subject to latency created by the stages, namely, (1) data acquisition from traffic participants (such as camera recordings and encoding, harvesting data from autonomous vehicles, and collect information from fiber); (2) transmission of data across communications links from sensors to inferencing servers using communications protocols such as Transmission Control Protocol (TCP), Unreliable Data Protocol (UDP) and Real-Time Streaming Protocol (RTSP); (3) data preprocessing (video decoding, and cropping); (4) AI inferencing; (5) higher level reasoning about required feedback to traffic participants; and (6) sending feedback to traffic participants across communications links via low-latency broadcast or dedicated channels over wired and wireless.
+3.4.1 Real-time requirements and low latency targets
+Sensor and control data in a real-time system is subject to latency created by the stages, namely, (1) data acquisition from traffic participants (such as camera recordings and encoding, harvesting data from autonomous vehicles, and collect information from fiber); (2) transmission of data across communications links from sensors to inferencing servers using communications protocols such as Transmission Control Protocol (TCP), Unreliable Data Protocol (UDP) and Real-Time Streaming Protocol (RTSP); (3) data preprocessing (video decoding, and cropping); (4) AI inferencing; (5) higher level reasoning about required feedback to traffic participants; and (6) sending feedback to traffic participants across communications links via low-latency broadcast or dedicated channels over wired and wireless.
 Smart city applications can be grouped according to their latency requirements.
@@
 This presents the opportunity to pursue novel engineering solutions and research problems.
-III-D2 Communication techniques and protocols
+3.4.2 Communication techniques and protocols
 Ultra-Reliable Low Latency Communications (URLLC), a key component of 5G wireless, can help achieve the low latency targets.
@@
 Tab.
-V summarizes the key characteristics of commonly used IoT communication protocols, Message Queuing Telemetry Transport (MQTT) [120], Constrained Application Protocol (CoAP) [121], and Hypertext Transfer Protocol (HTTP) [122].
+5 summarizes the key characteristics of commonly used IoT communication protocols, Message Queuing Telemetry Transport (MQTT) [161], Constrained Application Protocol (CoAP) [181], and Hypertext Transfer Protocol (HTTP) [60].
 Among them, MQTT emerges as a practical choice, due to its combination of low latency, high scalability, and reliable delivery mechanisms.
 Its lightweight publish/subscribe model, Quality of Service (QoS) guarantees, and session management make it well suited for real-time data exchange in dynamic, safety-critical environments like urban transportation systems [12, 20].
-TABLE V: Communication protocol comparison for real-time digital twin systems.
+Table 5: Communication protocol comparison for real-time DT systems.
 Protocol Latency Scalability Reliability Best Suited For
@@
 HTTP Higher latency due to new TCP/TLS handshake and headers per request Limited scalability due to client-server request/response model and verbose ASCII headers Relies on TCP for delivery, with no application-level acknowledgment, retries, or session handling Web APIs, periodic data transfer, backend integration
-IV Proposed DT pipeline
-Figure 2: Architecture of the proposed DT pipeline: Urban Transportation DT (UT-DT).
-In this section, we present a DT architecture for UT-DT, based on the sensing/communication/computing testbed deployed in NYC (Fig.
-2).
-The proposed architecture is enabled by cameras and LiDARs, high speed communications, and edge cloud computing.
-Figure 3: Diagram of intersection safety warning use case.
-IV-A Physical Infrastructure
-Pilot experiments are executed at the signalized intersection of Amsterdam Avenue (major) and 120th Street (minor) near Columbia campus in NYC.
-The road geometry and traffic statistics are summarized in Fig.
-4.
-Figure 4: Traffic statistics for NYC’s intersection of 120th St. & Amsterdam Ave.
-IV-A1 Technological enabler
-The physical functionality of the proposed UT-DT is built on top of the COSMOS testbed (“Cloud enhanced Open Software defined mobile wireless testbed for city-Scale deployment”) [125], developed for real-world research, development, and deployment of city-scale advanced wireless communications and innovative applications [125].
-It targets the technology “sweet spot” of ultra-high bandwidth and ultra-low latency, a capability that will enable a broad class of safety-time critical applications.
-Deployed in West Harlem, NYC, next to Columbia University campus, the COSMOS testbed is an enabler for research on the design, development, and deployment of a DT for urban traffic management.
-At the intersection of Amsterdam Avenue and 120th Street, cameras, LiDAR, and wireless sensing and communication nodes are deployed.
-The cameras employ H264 encoding with an I-frame interval set to 10 frames.
-The live RTSP video streams–with 4K resolution at 30 fps–were processed on a COSMOS edge server equipped with an A100 GPU.
-The LiDAR is a Velodyne VLP-32C with up to 200m range.
-IV-B Digital twin
-Building on data collected from the physical world, we built an Unreal Engine based simulation platform for generating photorealistic street scenes.
-Moving objects are populated into an integrated SUMO-CARLA simulation platform to validate our safety warning application [126], where vehicle locations are synchronized from the real-world MQTT messages.
-[…]
-(a) COSMOS in CARLA
-(b) COSMOS in SUMO
-(c) Boundless with snow
-(d) Boundless with rain
-Figure 5: (Top) Carla-SUMO simulator; (Bottom) Qualitative examples from Boundless. Bounding boxes for vehicles are shown in blue, and pedestrians are shown in green.
-IV-C Use cases
+4 Use cases for urban transportation
 The value of a DT for urban transportation management ranges from traffic state prediction, spatiotemporal traffic flow forecasting, to urban planning and policy making.
-Here we focus on two that are particularly important in urban settings, leveraging video analytics.
-IV-C1 Use case 1: Micromobility detection
-In the U.S., the e-bike had a market size of almost \$2 billion in 2022, with a projected growth of 15.6% from 2023 to 2030.
-In 2022, 1.1 million e-bikes were sold, four times as many as were sold in 2019.
-In 2019, 136 million trips were made using micromobility, a 60% increase from 2018.
-Because their presence and surge have profound implications for road safety and urban traffic management, automatic detection of micromobility for regulation becomes increasingly important.
-However, there does not exist a standard AI model nor a dataset for the object detection of this emerging transportation mode.
-AI model evaluation
-[…]
-TABLE VI: Class-wise recall (%) on the Micromobility test set. Higher values indicate better detection performance.
-Model Bicycle $\uparrow$ e-Bike $\uparrow$ Motorcycle $\uparrow$
-YOLOv9e 52.9 54.7
-Ours 68.1 61.0 91.8
-Figure 6: Four sample frames showing the small scale of micromobility instances from infrastructure-based 4K resolution cameras in the COSMOS testbed.
-IV-C2 Use case 2: VRU safety warning
-Safety-critical applications, making automated life-and-death decisions such as collision avoidance warning between automobiles and pedestrians, need to activate at the precise time and the right moment with bounded latency.
-Safety-critical systems are often time-critical and reactive, because they need to react to external signals in precise time [127], and require tasks to be executed in a timely manner.
+Here we primarily focus on those that are particularly important in urban settings leveraging video analytics, namely, VRU safety warning at road intersections and traffic signal control.
+4.1 Safety-time critical applications
+Safety-critical applications [6], making automated life-and-death decisions such as collision avoidance warning between automobiles and pedestrians, need to activate at the precise time and the right moment with bounded latency.
+Exemplary systems include autonomous driving, Industry 5.0, and mobile and collaborative robots, strictly entailing precise time, computing, communication with bounded latency and AI [23].
 Theses systems hold the substantial potential to enhance road safety and save lives.
@@
 The applications of DTs on safety-critical scenarios, however, remains understudied, because it necessities short runtime and real-time reaction, posing high requirements for communication and computing technologies, pipeline architecture design, and testing.
+Intersections, where sixty percent of crashes happen [88, 59], are critical bottlenecks of an urban transportation network.
+To improve urban road safety and increase traffic capacities, safety warning is the key.
 We summarize the safety-critical applications to address conflict risks between vehicles and VRUs in Tab.
-VII, and offer an outlook of safety guarantee related methods in Sec.
-V-A3.
-TABLE VII: Literature review on safety warning to VRUs.
+6, and offer an outlook of safety guarantee related methods in Sec.
+6.1.3.
+Table 6: Literature review on safety warning to VRUs.
 Ref Technology Objective Risk Assessment Metrics Evaluation Method/Metric
@@
 2. Distance error of vehicles and VRUs
-[126] Sensing: Camera, real time object detection
+[65] Sensing: Camera, real-time object detection
 Communication: LTE, Wi-Fi, and MQTT protocol
@@
 3. Distance to conflict point
-Ours Sensing: Camera, YOLOv8 object detection.
+[69] Sensing: Camera, YOLOv8 object detection.
 Communication&Networking: Fiber, LTE, and MQTT protocol
-Computing: Server end GPU
-Develops a DT pipeline to demonstrate use cases (including intersection safety warning and ATSC) in urban settings. TTC 1. Accuracy for the warning issued
+Computing: Server end GPU, resource allocation
+Develops a DT pipeline to demonstrate use cases (including intersection safety warning) in urban settings. Distance to the danger zone 1. Localization accuracy
 2. Granular latency per frame
-Intersections, where sixty percent of crashes happen [135, 136], are critical bottlenecks of an urban transportation network.
-To improve urban road safety and increase traffic capacities, safety warning is the key.
-Leveraging existing sensors at the intersection, we have developed a combined VRU and vehicle warning application.
-The cameras and the LiDAR sense the presence of VRUs at urban intersections, make predictions of their movements, apply traffic operation and control strategies, and feedback to system controllers and road users, with the primary goal of increasing traffic safety and efficiency (see Fig.
-3).
-Resource optimization
-Due to the computational needs of complex computer vision models and the high volume of video data, ensuring scalability in a video analytics pipeline requires optimization of its configuration to maintain performance.
-However, the system’s performance is impacted by factors such as the video content, network conditions, and available computational resources.
-As a result, it is crucial to implement a dynamic, real-time optimization mechanism that adjusts key configuration parameters—such as resolution, frame rate, and bitrate—based on these varying conditions.
-This adaptive approach allows the system to continuously balance performance with resource efficiency, ensuring scalable and reliable video analytics.
-Accordingly, we have equipped our video analytics pipeline with a missingResource Optimization (see Fig.
-2) component that continuously adapts the system’s configuration parameters to maintain its performance.
-In addition to dynamic adaptation of the system’s configuration, we can reduce latency and GPU consumption by identifying Regions of Interest (RoIs) where pedestrians might be in danger, using lightweight processing (e.g., low resolution, small models).
-As shown in Table VIII, smaller models (“YOLOv8s” in Column 4) incur significantly lower latency.
-Larger models (“YOLOv8x” in Column 6) with higher resolution and frame rate are only triggered when a critical danger area, i.e., RoI, is detected.
-For example, detecting large objects such as vehicles does not require large models or high resolution.
-Therefore, we can use a smaller model and lower resolution to detect vehicles and their trajectories, and based on that, determine the danger areas where pedestrians may be at risk or in the blind zones of vehicles.
-These identified danger areas are then processed using larger models and higher resolution to detect pedestrians at risk and notify them or the vehicles if necessary.
-In our intersection safety warning system (shown in Fig.
-3), we have embedded an RoI Specifier element to facilitate this approach.
-AI models
-A variety of AI models are developed to predict future trajectories of interacting road users at the intersection, including InfoSTGCAN (An Information-Maximizing Spatial-Temporal Graph Convolutional Attention Network) [137], which encodes road user trajectories into quantized latent codes to account for heterogeneity in road users, PI-NeuGODE (Physics-Informed Graph Neural Ordinary Differential Equations), using physics-informed graph neural ordinary differential equations [138, 139], and uncertainty quantification (UQ) to characterize the predictive confidence [140].
-These probabilistic trajectory prediction methods enhance pedestrian safety at intersections by incorporating UQ into risk assessment.
-The Kalman Filter (KF) [141] performs prediction by recursively estimating the state of a moving object using a linear dynamic model with Gaussian noise.
-Trajectron++ [142] models multimodal future trajectories by combining recurrent neural networks with conditional variational autoencoders, enabling probabilistic and socially-aware predictions.
-We present the predicted trajectories and performance comparison in Fig.
-7, where the performance is evaluated using Average Displacement Error (ADE, i.e., the mean Euclidean distance between predicted and ground-truth positions over all future time steps) and Final Displacement Error (FDE, i.e., the Euclidean distance at the final prediction step).
-Method ADE $\downarrow$ FDE $\downarrow$
-Kalman Filter 0.91 1.92
-Trajectron++ 0.77 1.34
-ADE and FDE comparison. Lower values are better ($\downarrow$).
-Figure 7: Performance comparison for trajectory prediction
-Evaluation
-In safety-critical applications, it is infeasible to perform field experiments for such assessment. DT simulation is thus used to assess the effectiveness of the system.
-a. End-to-end latency
-Measuring latency could be particularly challenging due to the need for cross-network timing synchronization of devices and compute server.
-Table VIII presents the latency incurred by the main components of a typical video analytics pipeline for object detection and tracking, as shown in Fig.
-3.
-We measured the latency for three sizes of YOLOv8 object detection model: YOLOv8s (small, $\sim$11.2 million parameters), YOLOv8m (medium, $\sim$25.9 million parameters), and YOLOv8x (large, $\sim$68.2 million parameters).
-Larger models offer improved detection accuracy but come at the cost of increased latency and higher GPU and memory usage [143].
-The pipeline has been systematically optimized in terms of memory and resource usage.
-These elements usually run on an edge device or server.
-The results identify potential bottlenecks within the pipeline and indicate which components could benefit from further optimization.
-TABLE VIII: Average latency and Standard Deviation for different pipeline elements.
-Metric Inference Elements Downstream Elements
-Reception Pre-processing Object Detection Model Object Tracking MQTT Msg Creator MQTT Msg Retrieval
-Small (YOLOv8s) Medium (YOLOv8m) Large (YOLOv8x) Ethernet Wi-Fi LTE 5G
-Avg Latency (Std Dev) [ms] 1.94 (1.69) 0.108 (0.024) 4.034 (0.084) 7.216 (0.086) 11.140 (1.800) 0.973 (0.173) 0.081 (0.021) 3.21 (0.315) 6.86 (1.19) 45.72 (15.30) 39.21 (7.12)
-b. Safety message generation accuracy
-Figure 8: A simulated pedestrian in CARLA (left) surrounded by cars with real-world positions (right).
-To validate the accuracy of our trajectory prediction and risk assessment algorithms for warning generation, we conducted three rounds of simulation, each lasting $10$ minutes and generating a total of $232$ virtual pedestrians in CARLA.
-We then compared the number of collision warning messages with the actual number of simulated collisions.
-The process is illustrated in Fig.
-8.
-(a) TTC (in seconds)
-(b) Distance (in pixels)
-Figure 9: ROC curves for trajectory prediction to determine optimal thresholds
-Note that metrics for risk assessment include time to collision (TTC) (i.e., the time remaining before a collision occurs) and post-encroachment time (PET) (i.e., the time interval between when the encroaching vehicle leaves the conflict point and when the vehicle of the right-of-way arrives at the conflict point).
-TTC is commonly used, while PET is typically for post-event analysis.
-To select the optimal thresholds of TTC, we first generate the ROC (Receiver Operating Characteristic) curve in Fig.
-9.
-Fig.
-9(a) illustrates the relationship between the true positive rate (TPR) and the false positive rate (FPR) across various TTC threshold values, ranging from 0.1 to 1.2 seconds.
-The optimal TTC threshold is 1.1 seconds (indicated by a blue square), at which the TPR reaches $0.958$ and the FPR is $0.4$.
-To compute TTC, we need to compare each pair of predicted trajectory points in the next $t$ time steps, determined by a danger distance threshold that is defined as when the proximity between a vehicle and a pedestrian constitutes a dangerous interaction.
-Fig.
-9(b) presents the TPR and FPR for the threshold of danger distances ranging from $5$ to $100$ pixels.
-The optimal threshold is valued at $30$ pixels, which yields a TPR of $0.958$ and an FPR of $0.345$.
-Based on the above two thresholds, we run the collision prediction model.
-The resultant confusion matrix for predicted collisions under the selected thresholds is given by $\text{TP},\text{FP}],[\text{FN},\text{TN}]]=[[66,45],[2,119]]$.
-c. Human response time assessment
-Figure 10: Response time determination in the VR experiment
-Figure 11: Survey results from the VR experiment
-Would issuing warnings to pedestrians help reduce users’ response time and increase their safety awareness?
-To test this hypothesis, we designed virtual reality (VR) experiments in Unity3D [144], where warnings are provided via voice and text displayed on the VR headset Meta Quest 3.
-There are two traffic scenarios, one involving an interaction between the participant (i.e., a pedestrian) and an oncoming scooter, and the other between the participant and an oncoming vehicle.
-In each scenario, participants first receive no warning and then a warning, and their response times are recorded.
-As illustrated in Fig.
-10, when traffic approaches at time $t$, the user either receives a warning or does not.
-The user stops after a response time $t_{r}$, which is then recorded.
-The response time distributions are presented in Fig.
-12, with the ‘no warning’ condition shown in yellow and the ‘warning’ condition in blue.
-The issuance of warning has reduced people’s average response time by $0.62\mathrm{s}$ (in Scenario 1) and $1.11\mathrm{s}$ (in Scenario 2), respectively.
-The demographics of the participants are shown in the upper section of Fig.
-11.
-The bar chart summarizes the participants’ survey responses regarding VR and traffic realism, the perceived importance of safety warnings, and comfort within the VR environment, with mean and standard deviation indicated on the top left corners of each subfigure.
-Figure 12: Comparison of response times with and without a warning. A scooter (in Scenario 1) and a vehicle (in Scenario 2) poses the danger, respectively.
-IV-C3 Discussions
-We would like to highlight that our NYC’s testbed deployment demonstrates generalizability.
-As the biggest metropolis of the U.S., NYC’s dense population, limited space, and multimodal transportation systems (consisting of public transit with buses and subway, driving, ferry) motivate us to study the interactions between automobiles and VRUs (including pedestrians, cyclists, and surging micromoblity and e-mobility).
-Located in one of the world’s busiest regions - Manhattan of the NYC, the COSMOS testbed offers a natural laboratory for us to collect abundant data and conduct live experiments, which would otherwise not be possible.
-We leverage the COSMOS testbed to design and implement sensor (e.g., camera, LiDAR) data analytics systems and evaluate them in urban settings to ensure they can scale to thousands of connected intersections.
-To achieve this, we ensure that the designed analytics systems are efficient in resource consumption, particularly network bandwidth and GPU usage.
-We also ensure that these systems are modular and deployable in a distributed manner across multiple edge servers and cloud infrastructure.
-Fig.
-3 demonstrates such a modular and distributable system.
-This design enables flexibility, allowing the system to be deployed in settings where a single edge server cannot handle the full workload, requiring distribution across several edge servers.
-Therefore, the system must be designed such that each module can be easily deployed on a different server and can seamlessly communicate with other modules, regardless of whether they reside on the same server.
-Additionally, the AI models used must be replaceable or generalizable to accommodate different environments, camera angles, heights, and other variations.
-With these principles in mind, we use the COSMOS testbed to develop scalable systems.
-The NYC’s dense high-rise buildings enable us to leverage legacy infrastructure for traffic sensing and monitoring, and facilitates the instrumentation of cameras and communication nodes.
-Note that there is a growing trend of studies relying on emerging communication protocols like C-V2X for vehicle DTs [145, 146], which requires customized infrastructure and user-end devices.
-For rapid and scalable deployment, our pilot experiments utilize contemporary protocols, which hold the potential to further leverage the city’s legacy traffic cameras and wireless communication infrastructure for scalability and cost-effectiveness.
-Our technologies could be replaced by emerging methods that can achieve lower latency and higher bandwidths.
-For instance, software defined features of the COSMOS testbed and the flexibility of radios make it possible to deploy emerging technologies such as C-V2X and experiment with novel low-latency applications.
-Since major intersections in the neighborhood has similar sensor suites, it would not need additional infrastructure deployment to generalize our system to network wide multi-intersection settings.
-We will extend our use cases to multi-pedestrian warning scenarios.
-Smartphone‑to‑cloud beaconing with an Adaptive Multi‑Mode (AMM) scheme enables real‑time location sharing from multiple devices, allowing the server to issue timely risk alerts [147].
-A Vehicle‑to‑Pedestrian (V2P) system [148] leveraging Bluetooth Low Energy (BLE) advertising on smartphones broadcasts standardized Personal Safety Messages (PSM) to address the multi‑device communication challenge in road safety.
-Moreover, efficient multi-device communication is essential for real-time data sharing and coordination among mobile users in smart urban environments.
-Key challenges include ensuring reliable connectivity in dense settings, minimizing communication latency, and maintaining user privacy.
-Advanced techniques such as device-to-device (D2D) communication [149], opportunistic networking [150], and privacy-preserving data aggregation [151] have been developed to address these issues.
-V Conclusions and open questions
+4.2 Adaptive traffic signal control
+Stochasticity arising from travel demands, un-predicted VRU movement, and traffic gridlocks requires traffic operators to design adaptive traffic signal control (ATSC).
+With a large amount of agents (including VRUs, traffic lights) continuously interacting in a stochastic environment, classical optimization tools in deterministic environments could fail to capture such complex decision-making processes.
+There is a surge in employing learning methods, including reinforcement learning and federated learning, to optimize traffic signals [95, 39, 104, 97, 264, 146, 112, 34, 79].
+With an increasing concern in data governance and privacy, federated learning witnesses a growing trend that allows optimizing centralized control while preserving distributed data privacy [243, 86, 15, 62, 56, 127].
+Related work are summarized in Tab.
+7.
+The last column indicates whether the optimization accounts for the safety of VRUs or not.
+Table 7: Summary of papers using (F)RL for ATSC
+Ref. RL FL DT Simulator Demand Network VRU
+[95] DDPG SUMO
+[39] SUMO
+[104] A2C Vissim
+[97] Vissim
+[264] SUMO
+[62] A3C FedAvg SUMO
+[243] DQN FedAvg Cityflow
+[86] PPO FedAvg SUMO
+[15] DQN FedAvg SUMO
+[146] A2C SUMO
+[112] A2C SUMO
+[34] A2C SUMO
+[79] PPO SUMO
+[167] PPO SUMO
+[173] D3QN SUMO
+[241] DDQN SUMO
+Demand: Real-world traffic demand data; Network: Real-world road network.
+FedAvg: Federated Averaging, DDPG: Deep Deterministic Policy Gradient, A2C: Advantage Actor Critic, A3C: Asynchronous Advantage Actor Critic, PPO: Proximal Policy Optimization, DQN: Deep Q Learning.
+5 Physical Testbeds for DT validation
+Physical testbeds are essential for verification, validation, and testing of DTs in diverse and out-of-distribution scenarios.
+It offers a platform for sensor instrumentation, data collection, actuation, to close the information flow and control feedback loop.
+Datasets collected from a testbed, if open-sourced, could be used by other researchers who cannot afford building a testbed.
+Below, we summarize existing testbeds in Tab.
+8 based on their types, namely, real-world, closed-track, or scaled-down.
+Except COSMOS focusing on the dense urban traffic environment, these testbeds are primarily developed for the testing of vehicular technologies and control.
+Table 8: Summary of real-world and scaled down testbeds in the U.S.
+Type Testbed Open Data Software Hardware Testing scenarios
+Real-world COSMOS (NY) [171] software-defined networking (SDN), edge cloud, radio software stack radio access & wireless nodes, mmWave antennas, cameras, GPU/FPGA accelerators V2X networking; traffic sensing; edge computing experiments
+Toyota Testbed (CA) [215] AWS cloud platform, traffic simulation platform edge computing, vehicle on-board sensors, full-scale vehicle, onboard computing units, CPU/GPU nodes CAV platooning, ATSC algorithm evaluation, mixed traffic, vehicle control testing
+SMTG (NJ) [92] computing cluster, Carla/SUMO co-simulation, Unity visualization lidar, camera, Bluetooth, NVIDIA Jetson AGX Xavier on edge, RSU, Cellular modems real world traffic detection, V2X, DT evaluation, VRU safety.
+STARlab (WA) [210] STAR Detection App, VISSIM simulation platform, edge station software V2X Edge Node, Bluetooth and WiFi radios, Raspberry Pi V2X safety, mobility detection
+GoMentum (CA) [35] HD Map Database, V2X middleware, sensor middleware full-scale AVs, edge computing, lidar camera, RTK-GPS signalized intersection testing, Construction zone testing, VRU testing
+Purdue Testbed (IN) [263] autonomy stack, ROS, cloud-based VLM, whisper, Autoware full-scale AV, lidar, radar, camera, GNSS, high bandwidth networking highway driving, intersection safety, VLM-controlled AV, parking lot testing
+Closed track Mcity (MI) [163] unreal virtual Mcity, V2X software stack, driving simulation 32-acre closed test track, perception sensors, RSUs, DSRC, Roadside LiDAR, full-scale AVs ATSC, CAV testing, signalized intersections, AV perception/decision experiments
+Scaled down Duckietown (MA) [162] Robot Operating System (ROS) server, signal control framework scaled autonomous vehicle and road, traffic light, Wi-Fi lane following, city navigation, traffic light control
+F1Tenth (MA,PA,VA) [14] ROS, Gazebo simulation, RViz visualization, oracle node 1/10 autonomous racecar, NVIDIA Jetson TX2, Wi-Fi, UDP sim-to-real, vehicle racing, vehicle control
+Ccity (NY) [168] ROS, self-built simulation, driving emulator 1/8 vehicle with Jetson Orin Nano, camera, UWB localization, traffic signal, Wi-Fi, Raspberry Pi AV lane following, platooning, AV control, traffic signal control, human driving behaviour testing
+IDS3C (NY) [26] UDP sockets, IDS 3D city, ROS, Unity engine, Microsoft AirSim plugin 1/25 robotic ground cars, Raspberry Pi, Wi-Fi, UDP, VICON motion capture system CAV control, CPS validation, eco driving, shared mobility, sim-to-real
+6 Conclusions and open questions
 In this paper, we first review the AI methods applied to every stage of the DT pipeline, from object detection, tracking, prediction, simulation, to traffic operation and management.
-Then, leveraging the unique characteristics of NYC, we propose a DT architecture and present a safety-critical use case, namely, intersection safety warning.
-Three evaluation methods are performed, including measuring latency, simulating collision reduction, and VR experiments for human responses.
-A growing number of sensors, explosive amounts of data, and increasing computational powers have opened up tremendous opportunities for researchers to apply AI to create, train, evaluate, and streamline DT pipelines for uban traffic management.
+Then, how DTs are leveraged for use cases including intersection safety warning and traffic signal control are introduced.
+Following that, the testbeds for DT testing and validation are presented.
+A growing number of sensors, explosive amounts of data, and increasing computational powers have opened up tremendous opportunities for researchers to apply AI to create, train, evaluate, and streamline DT pipelines for urban traffic management.
 Subsequently, we will present emerging trends, challenges, and open questions for the development of DT.
-V-A Emerging trends
+6.1 Emerging trends
 While literature on DTs for individual components has been surging, how each element of the DT works collectively and function organically is key to the development of the next-generation DT, which could empower the intelligence and automation of transportation applications.
-V-A1 Engineering the pipeline
+6.1.1 Engineering the pipeline
 Engineering a DT via the integration of multiple subsystems poses technical challenges, and we will name a few.
@@
 Future research directions may focus on leveraging high angular resolution from densely packed phased arrays to achieve precise beamforming, with the potential to introduce imaging-like capabilities within communication networks [159, 76].
-V-A2 Emerging DT applications
-With the increasing demand for urban passenger and goods delivery, emerging technology like urban sensing, electrification, connectivity and autonomy, and robotics have been gradually transforming urban streetscapes, which pose new challenges to the operation and management of urban infrastructure and public space.
+6.1.2 Emerging DT applications
+With the increasing demand for urban passenger and goods delivery, emerging technologies like urban sensing, electrification, connectivity and autonomy, and robotics have been gradually transforming urban streetscapes, which pose new challenges to the operation and management of urban infrastructure and public space.
 DTs are crucial to improve urban safety and mobility in infrastructure planning, service operation and management, and ultimately, policy making.
@@
 In Fig.
-13, emerging use cases are categorized based on required communication latency (x-axis), spatial resolution (y-axis), and data bandwidth (z-axis), respectively.
+2, emerging use cases are categorized based on required communication latency (x-axis), spatial resolution (y-axis), and data bandwidth (z-axis), respectively.
 Figure 2: Urban T-DT use cases.
-Stochasticity arising from travel demands, un-predicted VRU movement, and traffic gridlocks requires traffic operators to design adaptive traffic signal control (ATSC).
-With a large amount of agents (including VRUs, traffic lights) continuously interacting in a stochastic environment, classical optimization tools in deterministic environments could fail to capture such complex decision-making processes.
-There is a surge in employing learning methods, including reinforcement learning and federated learning, to optimize traffic signals [169, 170, 171, 172, 173, 174, 175, 176, 177].
-With an increasing concern in data governance and privacy, federated learning witnesses a growing trend that allows optimizing centralized control while preserving distributed data privacy [178, 179, 180, 164, 181, 182].
-Driven by growing urban populations, low-altitude economy aims to exploit vertical space in urban environments, advanced by drones and eVTOL (electric vertical take-off and landing).
-It will foster wide applications, ranging from urban air mobility, urban logistics, agriculture, emergency service, to infrastructure inspection, surveillance and security.
-Critical challenges arise in sensing, modeling, predicting urban aerodynamics, and collision avoidance with high-rise buildings and flying objects.
-Open questions include market demand estimation, infrastructure planning, operation and design.
-DTs hold the potential to the adoption of these emerging technologies, for its power in simulating emerging demand patterns, modeling their kinetics and dynamics, routing and charging behaviors.
-To build such DTs, CPS technologies and AI methodologies would facilitate precise sensing and perception [146], reliable and fast communication and networking [183], seamless multimodal coordination (such as drone-truck-mobile robot delivery) [184], as well as safe and optimal decisions [185].
-V-A3 AI models
+6.1.3 AI models
 AI methods have been applied to every stage of the DT pipeline. We will point out key challenges of AI methods in transportation.
@@
 DTs play important roles in the above tests, while experimental design is crucial to cover comprehensive scenarios leveraging game theory [145] and in recent years, generative models [147].
-The presence of human factors along the DT pipeline could complicates the evaluation of cyber-physical-human systems, because of randomness and unpredictability in human behaviors [208].
+The presence of human factors along the DT pipeline could complicate the evaluation of cyber-physical-human systems, because of randomness and unpredictability in human behaviors [42].
 Humans are not only participants in traffic (as drivers, pedestrians, cyclists), they are also creators and users of DTs.
@@
 Since there does not exist a unified approach for the DT testing, evaluation methods, metrics, testing platforms should be developed to facilitate standardized assessment and validation of AI models.
-Benchmark datasets and methods
+Benchmark datasets & shared testbed resources
+Due to limited availability and single perspective of real-world datasets, synthetic data generation from realistic 3D simulators has become increasingly popular for urban computer vision problems.
+However, open mixed-perspective datasets that adapt to diverse deployment conditions remain scarce.
+Advances in open-vocabulary object detection offer a promising avenue to address this gap [129, 144, 30].
 To advance the application of AI in DT, we must “stand on the shoulders of giants.” In other words, each team does not simply develop one DT for its own internal use.
@@
 Thus, the transportation community must push to open source data, codes, simulation, algorithms, and results for replicability.
-For example, standard test platforms, such as Gym [214], Flower [215], PettingZoo [216, 217], and Ray RLlib [218], are important to benchmark various AI models and methods, which is mostly missing in the transportation community.
-Moreover, open mixed-perspective datasets that adapt to diverse deployment conditions remain scarce.
-Advances in open-vocabulary object detection offer a promising avenue to address this gap [219, 220, 221].
-V-B Challenges and open questions
-V-B1 Closing real-to-sim-to-real gap
+For example, standard test platforms, such as Gym, Flower, PettingZoo, and Ray RLlib, are important to benchmark various AI models and methods, which is mostly missing in the transportation community.
+While data and methods can often be shared, advancing DTs ultimately depends on shared physical testbeds.
+These real-world platforms are essential for validation and learning, yet remain costly and space-constrained—calling for new, collaborative approaches from the community.
+Recent years have seen a growing effort to crowd-source existing testbeds pushed by both agencies and researchers.
+For example, augmented from the Mcity Test Facility by remote access, Mcity 2.0 [150] enables researchers to connect to the site via a secure cloud interface, run experiments without being physically present, and interact with a mixed-reality environment that overlays virtual actors onto the physical track.
+In this way, Mcity 2.0 operationalizes standardized, repeatable testing by tightly integrating simulation, physical infrastructure, and remote vehicle control in one unified platform.
+6.2 Challenges and open questions
+6.2.1 Closing real-to-sim-to-real gap
 A key challenge in applying models trained in simulated environments to the real-world is domain shift between the training and test environments.
@@
 Such a shift – caused by discrepancies between simulated and real-world conditions – can undermine the generalization of models, resulting in reduced performance in scenarios not represented in training.
-What distinguishes an DT from a conventional computer simulator?
-We believe the key lies in its capability of generating simulations and predictions consistent with the real world, as well as preserving intervention performance.
 Accordingly, two gaps exist while establishing a DT, namely, real-to-sim (real2sim) gap (i.e., the deviation of the simulated digital world from the real-world), and sim-to-real (sim2real) gap (i.e., the performance deviation of interventions implemented in the real-world from those simulated in the digital world).
@@
 Although these gaps penetrate through each stage along the DT pipeline, from computer vision, tracking, to prediction, and intervention, sim2real transfer is more studied in object detection [165] and policy learning [164, 178, 134, 36].
-The major application area of sim2real transfer in DT is robotic manipulation [227], while that in urban navigation and autonomous driving has witnessed a gradual surge, participially in computer vision [228] and reinforcement learning [229, 230].
+The major application area of sim2real transfer in DT is robotic manipulation [1], while that in urban navigation and autonomous driving has witnessed a gradual surge, particularly in computer vision [229] and reinforcement learning [83, 109].
 Key methods include domain randomization [193, 147], which introduces variations to augment training data and expose the model to a wide range of scenarios; and domain adaptation [195, 227], a technique of finding mappings to transfer data points observed empirically from two different data distributions.
@@
 Domain knowledge could help represent the physical world with higher accuracy.
-For example, hybrid twin [235, 236] relies on both data-driven and physics-informed AI [237].
+For example, hybrid twin [33] relies on both data-driven and physics-informed AI.
 Since calibration of a full model could be time-consuming, expensive, and potentially infeasible, the model reduction philosophy has become popular.
@@
 Digital cousin [37] aims not to build a simulation model that replicates the reality exactly, instead, mainly focus on end-to-end gap, namely, from real-world sensing to intervention.
-V-B2 Prototyping DTs for human intelligence
+6.2.2 Prototyping DTs for human intelligence
 Many studies define a DT as emerging technologies, namely, object detection and tracking with edge-cloud computing and communication.
@@
 Thus, we believe that GenAI-powered DT will be the next-generation of DTs for efficient and safe traffic management.
-V-B3 Limits in AI
+6.2.3 Limits in AI
 Despite the promising future of AI-powered DTs, application of AI to DTs could face challenges.
@@
 We hope this paper can serve as a pointer to help researchers and practitioners understand SOTA methods and gaps on the development of DTs; a bridge to initiate conversations across interdisciplinary researchers; and a road map to exploiting potentials of DTs for urban transportation applications.
+Acknowledgments
+This work was supported by NSF CPS-2038984 and ERC-2133516. We thank Mengxuan Liu for her assistance in generating the 3D models.